### v1

# Real-Time Position-Aware View Synthesis from Single-View Input

Abstract.
Recent advancements in view synthesis have significantly enhanced immersive experiences across various computer graphics and multimedia applications, including telepresence, and entertainment. By enabling the generation of new perspectives from a single input view, view synthesis allows users to better perceive and interact with their environment. However, many state-of-the-art methods, while achieving high visual quality, face limitations in real-time performance, which makes them less suitable for live applications where low latency is critical. In this paper, we present a lightweight, position-aware network designed for real-time view synthesis from a single input image and a target camera pose. The proposed framework consists of a Position Aware Embedding, modeled with a multi-layer perceptron, which efficiently maps positional information from the target pose to generate high dimensional feature maps. These feature maps, along with the input image, are fed into a Rendering Network that merges features from dual encoder branches to resolve both high level semantics and low level details, producing a realistic new view of the scene. Experimental results demonstrate that our method achieves superior efficiency and visual quality compared to existing approaches, particularly in handling complex translational movements without explicit geometric operations like warping. This work marks a step toward enabling real-time view synthesis from a single image for live and interactive applications.
1. Introduction
Telepresence systems (Dima and Sjöström, 2021) enable users to interact with remote environments in real time, offering a wide range of applications, particularly in sectors such as heavy machinery, where remote operation from a safe distance is crucial. These systems have the potential to facilitate seamless natural interaction with remote environments. Current telepresence technologies typically rely on head-mounted displays (HMDs), high-speed networks, and consumer or industrial-grade cameras to deliver live video feeds of the environment. While 2D video feeds provide some pictorial information, the lack of depth perception significantly limits situational awareness, which is vital for effective remote operation.
The challenge to have an understanding of a remote site’s structure may be address by enabling motion parallax for the user. I.e., the user moves the head forth and back to get a notion of the scene depth. This requires access to views from different perspectives, where camera positions and poses need to be continuous for a smooth transition and a high quality experience. Therefore, it is imperative to render continuous perspective views accurately and in real time for applications such as augmented reality (AR), virtual reality (VR), and telepresence systems.
The synthesis of views from perspectives continuous in camera position and pose requires sufficient information of the scene. Early methods employed view interpolations and depth-image-based rendering, but both these require multiple capture devices resulting in calibration issues and high communication capacity requirements. Therefore, generating new views from a single input image and a specified camera pose has emerged as a key task in rendering. A variety of approaches, such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and more recent transformer-based architectures like NViST (Jang and Agapito, 2024), have been proposed to tackle this problem. These methods leverage both synthetic and real-world datasets to produce high-quality novel views. However, most state-of-the-art methods either require multi-view inputs or are computationally intensive, making them unsuitable for real-time applications where low latency is critical.
In this work, we build upon the growing body of research on single-image-based view synthesis. We introduce a learning based position embedding method combined with a rendering network that efficiently synthesizes high-quality views of arbitrary camera location and pose using fewer computational resources. Our method maintains high fidelity even when handling large translational shifts, making it ideal for real-time applications and practical deployment on consumer devices. To train and test our method we also create our own synthetic dataset using Blender.
Unlike other methods, such as NViST, which employs transformer architectures, or SE3D (Koh et al., 2023), which utilizes large-scale generative-adversarial networks (GANs) to address occlusions and complex scenes, our network is designed to be lightweight and computationally efficient. Despite its smaller footprint, it excels in handling scenarios where translational movements dominate, enabling real-time inference. In summary, our main contributions are as follows:
A lightweight network with a positional embedding module that can generate novel views within a given baseline without requiring explicit operations.
A positional encoding scheme similar to NeRF, which is combined with an multi-layer perceptron (MLP), to create a more generalizable framework capable of handling different scenes.
Our method achieves real-time performance, with inference times of approximately 8.6 milliseconds, corresponding to over 100 frames per second (FPS) for images at a resolution of 512x512 pixels.
The remainder of this paper is organized as follows: Section 2 reviews state-of-the-art methods aimed at view synthesis. Section 3, and 4 describes the proposed approach for small baseline and large baseline rendering. Section 5 provides a comparative analysis of the results, including quality and speed evaluations. Finally, conclusions are summarized in Section 7.
2. Background
View synthesis has recently become a pivotal research area in immersive telepresence, aiming to generate unseen perspectives of a remote scene from a limited set of input images. Traditional approaches, such as structure-from-motion (SfM) and multi-view stereo (MVS), rely on geometric consistency between input views to reconstruct novel viewpoints (Seitz and Dyer, 1999). While these methods have shown effectiveness in controlled environments, their performance deteriorates in complex real-world scenarios and often necessitates multiple input views to achieve high-quality results. In light of these limitations, more recent methods focus on synthesizing novel views from a sparse or single image. This section outlines these advancements and their respective contributions to the field of view synthesis.
2.1. Scene Regression Approaches
Recent advancements in view synthesis have centered around methods that aim to represent a scene from sparse image inputs, particularly those leveraging neural implicit representations such as NeRF (Mildenhall et al., 2021). A NeRF encodes a scene as a continuous volumetric function, allowing novel view generation through volumetric rendering. Despite its capacity to produce highly photorealistic images, NeRF suffers from computational inefficiencies, particularly in terms of inference time. Several techniques have been proposed to mitigate these limitations, such as hash encoding in InstantNGP (Müller et al., 2022) and multi-scale representations in Mip-NeRF (Barron et al., 2021). While these improvements accelerate rendering, they retain significant reliance on MLPs, which remain computationally expensive.
To address these concerns, methods like Plenoxels (Fridovich-Keil et al., 2022) have emerged, leveraging spherical harmonics to represent a scene without the need for deep neural networks. Though Plenoxels achieves faster optimization than many NeRF-based methods, it still falls short in terms of memory efficiency. A more recent innovation, Gaussian Splatting (Kerbl et al., 2023), improves memory and rendering efficiency by employing high-dimensional Gaussian kernels to represent the scene. However, its optimal performance still depends on the availability of multiple input views.
2.2. Light Field Reconstruction and Rendering
Light field (LF) representation encodes both the spatial and angular information of a scene, often described by an HxWxNxN structure, where HxW refers to the spatial resolution and NxN denotes the angular resolution (Levoy and Hanrahan, 1996). This representation facilitates tasks like viewpoint manipulation (Zhou et al., 2020). LF reconstruction methods aim to generate a dense NxN light field from a single input view or a few input views, eliminating the need for complex multi-camera setups.
A common LF reconstruction approach utilizes convolutional neural networks (CNNs) to estimate scene depth (Srinivasan et al., 2017), followed by warping operators to generate target views. To address warping artifacts, subsequent refinement networks have been developed (Zhou et al., 2020, 2021). Other methods exploit depth estimation or optical flow (Cun et al., 2019), appearance flow (Ivan et al., 2019), or a combination of both (Bae et al., 2021) to synthesize the light field.
In more recent developments, GANs (Chen et al., 2020; Chandramouli et al., 2020) and auto-encoders (Han and Xiang, 2022) have demonstrated effectiveness in reconstructing dense light fields from a single image input. Additionally, LFSphereNet (Gond et al., 2023) extends LF reconstruction to spherical light fields using a single omnidirectional image. Furthermore, to generate the novel views from the LFs, the algorithms like dynamically reparameterized LFs (Isaksen et al., 2000) and per-view disk-based blending (Overbeck et al., 2018) have to be utilized which have higher memory footprint, making the whole process unsuitable for dynamic environments and low memory systems.
Our proposed method, PLFNet (described in Sec 3), contributes to LF reconstruction and rendering by directly synthesizing views within a local LF volume, thus eliminating the need of additional rendering algorithms.
2.3. Learning-Based Approaches
Learning-based methods seek to overcome the limitations of both scene regression and light field rendering approaches, which include constraints such as low baselines, sparse image input requirements, and slow rendering times. Several models, including PixelNeRF (Yu et al., 2021) and VisionNeRF (Lin et al., 2023), have been developed to generalize NeRF to novel scenes by conditioning the network on image features. Despite their promise, these methods often face challenges such as scale ambiguity and difficulty in handling large translational movements.
To address these challenges, the diffusion model based approach in 3DiM (Watson et al., 2022) employs pose guided diffusion to improve performance, particularly in scenarios with large translational movement. Fine-tuning diffusion models on large-scale synthetic datasets (Deitke et al., 2023) has also shown potential (Liu et al., 2023, 2024), although these methods are hindered by slow inference times.
An alternative strategy involves using multiplane image (MPI) representations, which approximate a scene using a series of parallel planes at fixed depths, each containing an image and alpha map. This representation allows novel view generation based on a fixed (Zhou et al., 2018; Flynn et al., 2019; Tucker and Snavely, 2020; Li and Kalantari, 2020) or variable (Li and Kalantari, 2020) number of planes. While MPI-based methods achieve superior visual quality, their computational demands remain high, especially during MPI layer prediction. Large-scale GANs (Koh et al., 2023) have also been explored for their ability to render fast and handle large baselines, though they introduce significant memory overhead, limiting their practicality.
More recent innovations, such as NViST (Jang and Agapito, 2024), incorporate vision transformers (ViTs) (Dosovitskiy, 2020) into view synthesis architectures, demonstrating strong generalization capabilities in real-world settings. By leveraging masked autoencoders (MAEs) and adaptive layer normalization conditioned on camera parameters, NViST delivers high-quality results, even for out-of-distribution scenes.
Our proposed method, PLFNet+ (described in Sec 4), contributes to the learning-based category by extending the views from a single input image while diverging from diffusion and transformer-based methods. By utilizing a lightweight design, PLFNet+ reduces computational overhead, enabling real-time rendering. This efficiency makes it particularly suited for dynamic environments where fast processing is crucial, addressing key challenges in the field of view synthesis.
3. Method: Light Field Reconstruction - PLFNet
We first define a Position Aware Light Field Reconstruction network called ”PLFNet” which aims to reconstruct the LF from a single input image by iterative reconstructing individual sub-aperture images. The source image serves as the basis for reconstructing the surrounding LF , where x represents the angular coordinate , and signifies the center of the LF. Each sub-aperture image at location , specified as , undergoes reconstruction via:
This two stage pipeline has been shown in Figure 1
3.1. Positional Embedding
We define the target position by the coordinate vector , where , and represent translations associated with angular coordinates. Given limited baseline shifts of LF reconstruction, directly using normalized values of can suffice for training the network on discrete locations but querying for any continuous values of .
Coordinate Normalization: The target coordinates are normalized into to ensure consistency across varying ranges of , . The normalization process is given by
This normalization step ensures that the coordinates lie within a fixed range, which is crucial for subsequent encoding.
Input Channel: To incorporate the normalized coordinates into the network, the values of are expanded into a tensor , where the first channel holds the values, and the second channel holds the values
This tensor has the same spatial resolution as the input image, ensuring that positional information is aligned with the image features spatially.
3.2. Rendering Network
The role of the rendering network in PLFNet is to synthesize a novel target image from the input image and the positional embedding , obtained from the positional embedding stage. Our network architecture is inspired by LFSphereNet (Gond et al., 2023) and consists of two distinct encoder branches followed by a decoder network. These branches work in tandem to extract and merge spatial and positional features, ultimately generating high quality novel views.
The rendering network operates as follows: the first encoder processes the input image to extract image-based features, while the second encoder processes both the input image and the positional feature map to capture position-dependent information. These features are subsequently merged in the decoder, which refines and upsamples the combined feature maps to generate the final output , a novel view at the target position.
3.2.1. Encoder I - Pre-trained Image Feature Extractor
The first encoder, , is a pre-trained image feature extractor that processes the input image . Following the approach used in ResNet-152 (He et al., 2016), we utilize the initial layers of ResNet to extract lower-level feature maps that are rich in texture and local detail. Specifically, extracts a set of feature maps , where and represent the spatial dimensions of the downsampled feature maps. These feature maps are then expanded to through additional convolution layers to increase the depth of the feature space which can be described as
The extracted features capture essential image content but lack positional awareness, which is handled by the second encoder.
3.2.2. Encoder II - Position-Aware Feature Extractor
The second encoder, , integrates both the input image and the positional embedding to adaptively capture position-specific features. Unlike the first encoder, which only processes the image, focuses on refining the feature maps based on the target position.
We first concatenate the input image and the positional embedding along the channel dimension, resulting in a stacked input tensor
where is the number of channels in . In our case, for LF Reconstruction.
This stacked tensor is passed through a series of convolutional layers that preserve the spatial dimensions, ensuring pixel-level positional information is retained throughout the encoding process, as demonstrated by LFSphereNet (Gond et al., 2023). The encoder produces a feature map , where is the number of feature channels, and skip connections are generated from intermediate layers for later use in the decoder.
Formally, this can be expressed as
The skip connections, denoted as , are collected from intermediate layers, where indexes the corresponding layer, and are used to enhance spatial detail during the decoding stage.
3.2.3. View Decoder
The decoder, , is responsible for merging the feature maps from both encoder branches and reconstructing the target image at the desired target position. The concatenated feature maps and are first combined along the channel dimension
This concatenated feature map is then passed through a series of transposed convolution layers to progressively upsample the spatial resolution, ultimately matching the resolution of the input image. The decoder uses a combination of transposed convolutions and bilinear upsampling, preserving spatial coherence in the generated image. Skip connections from the second encoder are concatenated into the upsampled feature maps at various stages to retain high-frequency details. The skip connections ensure that fine-grained positional and image details are retained, resulting in more consistent and accurate novel views. The overall process can be expressed as
The final output of the decoder is the synthesized target image , with spatial dimensions .
3.3. Training Loss
To optimize our PLFNet, we employ a combination of the L1 loss and the Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003) loss. This combination leverages the complementary strengths of each loss function. The L1 loss minimizes pixel-wise differences between the predicted sub-aperture images and the ground truth images , which helps to maintain accurate reconstruction of intensity values. The MS-SSIM loss, on the other hand, is a perceptually motivated metric that focuses on preserving structural information and texture across multiple scales, which correlates better with human visual perception as shown in image restoration task (Zhao et al., 2016).
The total loss function is defined as a weighted sum of the L1 and MS-SSIM losses
where is a weighting factor that balances the contribution of the two losses. This combined loss function enables the network to generate high-quality light field reconstructions that are both pixel-accurate and perceptually consistent.
3.4. Implementation: PLFNet
The input of PLFNet is a single RGB image and the target position . Here and which are similar to the spatial resolution of input image . We set because the datasets (Srinivasan et al., 2017; Raj et al., 2016; Rerabek and Ebrahimi, 2016) are organized as a 2D grid, with only two angular coordinates considered. PLFNet was implemented in PyTorch and trained with the following hyperparameters: batch size of 24, learning rate of 0.003, adam optimizer, MS-SSIM and L1 loss. After each 30 epochs, the learning rate was decreased by a factor of 0.2. The network was trained for 150 epochs on 2 GPUs (Nvidia A40) within a computing cluster with Intel Xeon Gold 6338 CPUs. The training duration was 48 hrs. For the dataset split, 75% of the data was used for training, 12.5% for validation and 12.5% for testing.
4. Method: View Synthesis - PLFNet+
Our approach for view synthesis builds upon PLFNet by extending the positional embedding to handle wider baselines with translation along the -axis. Here, represents the synthesized image corresponding to the target viewpoint. Our approach consists of two key stages as shown in Figure 2: first, the target position is embedded into a high-dimensional space through positional embedding (discussed in Section 4.1), and second, novel views are rendered from these positions, which is detailed in Section 4.2.
4.1. Extended Positional Embedding
We define the target position by the coordinate vector , where , , and represent translations from the input image . As the baseline increases and the search space extends along the axis, it becomes necessary to project the target coordinates into a higher-dimensional space to accurately capture positional information. To address this, we employ a positional embedding strategy that combines positional encoding, inspired by NeRF (Mildenhall et al., 2021), along with a MLP for enhanced representation of the target coordinate .
Coordinate Normalization: Similar to the normalization strategy in Section 3.1, we convert the target coordinates into their normalized form .
Positional Encoding: To project the normalized coordinates into a higher-dimensional space, we apply positional encoding. This involves concatenating multiple sine and cosine functions of varying frequencies
where is the normalized 3D coordinate vector, represents the encoding depth, and . The parameter is a scaling constant based on the domain of . For our experiments, we use , therefore yields an output of size dimensions, which provides enhanced positional awareness (see Table 6 for ablation study results).
Learning Based Reprojection: Once the positional encoding has been computed, we further enhance the representation by passing it through an MLP. This MLP maps the low-dimensional encoded vector into a higher-dimensional space, and it has proven to work well with scene regression task as shown in NeRF (Mildenhall et al., 2021). This process allows for a more expressive and flexible feature space, and our approach is similar to the adaptive layer normalization techniques used in NViST (Jang and Agapito, 2024).
Let be the output of the positional encoding, where is the dimensionality of the encoding. The MLP projects this vector into a high-dimensional space , where is chosen based on the desired resolution. Mathematically, the projection is represented as
After this, the vector is reshaped into a feature map of size , where . This reshaping operation can be expressed using the inverse vectorization notation as
Finally, the feature map is upsampled to match the spatial resolution of the input image, denoted by . This upsampling is performed using a scaling factor , as follows
The final upsampled representation retains the positional information of the target coordinate and is subsequently used by the rendering network to generate the novel view.
4.2. Rendering Network
The rendering network remains similar to Section 3.2. However, in the case of view synthesis, the positional embedding is denoted by a single channel, simplifying the input to the second encoder as
The output of the rendering network is the target image .
4.3. Training Loss
For the view synthesis task, we extend the loss function previously defined in Section 3.3 to better capture fine-grained details, especially with larger baselines. We maintain the same combination of L1 loss and MS-SSIM loss, which focuses on minimizing pixel-wise differences and preserving structural consistency, respectively. However, to further enhance the quality of the synthesized views, particularly in the frequency domain, we incorporate the Focal Frequency Loss (FFL) (Jiang et al., 2021) into the total loss function.
The FFL addresses the spectral bias in neural networks by adaptively focusing on the frequency components that are hardest to synthesize. This allows the model to improve the reconstruction of high-frequency details, which are critical for realistic novel view generation.
The FFL is defined as
where and represent the real and predicted frequency components at position in the Fourier domain, and is a dynamic spectrum weight that emphasizes hard-to-synthesize frequencies.
Thus, the total loss function for PLFNet+ is updated to
where is the Focal Frequency Loss, and is a weight balancing its contribution relative to the L1 and MS-SSIM losses.
This augmented loss function allows PLFNet+ to generate sharper, more realistic novel views by not only optimizing for spatial accuracy but also addressing frequency domain discrepancies.
4.4. Implementation: PLFNet+
We keep the same implementation as in PLFNet (described in Section 3.4, but set , . We also set as the positional maps is single channel. Due to limitation of training data when using our synthetic dataset we employ multi stage training strategies to reach satisfactory results in training process.
4.4.1. Multi Stage Training
Given the scarcity of large-scale datasets for view synthesis with wide baselines, we adopt a three-stage training approach shown in Figure 3 to progressively enhance the network’s ability to capture image features and handle positional shifts.
Stage 1 serves as a preconditioning step, training the network on a large-scale light field dataset (Srinivasan et al., 2017). While the baseline between the views in each light field is relatively narrow, the dataset’s large scale enables the network to learn both high- and low-level image features, implicitly understanding depth. In this stage, the network is conditioned to synthesize novel views within a constrained light field volume, similar to the reconstruction process described in PLFNet (Section 3). Once the network demonstrates proficiency in this task, we transition to the next stage.
Stage 2 introduces a dataset with a significantly larger baseline: the Spherical Light Field Database (SLFDB) (Zerman et al., 2024), which consists of 20 spherical light fields captured with 60 angular dimensions. From these omnidirectional images (ODIs), we extract planar patches and fine-tune the network with a lower learning rate. This step enables the network to learn to handle wider baseline movements while maintaining positional accuracy. Following this, the network is prepared for the final stage of training, which focuses on the joint optimization of the MLP and rendering network with more random positions.
Stage 3 involves training the network on a synthetic dataset containing random camera viewpoints within an cubical volume. The increase in the spatial search space necessitates the use of the positional embedding module for accurate view synthesis. In this stage, the network is fine-tuned to handle more complex camera translations and larger scene variations, fully utilizing the positional embedding to enhance rendering precision.
5. Experimental Evaluation
In this section, we present a comprehensive evaluation of our proposed PLFNet and PLFNet+ architectures. We begin by assessing the performance of PLFNet+ for view synthesis on datasets with varying baselines and complexities. Following this, we demonstrate the capability of PLFNet in reconstructing full light fields from a single input image. The experiments for view synthesis are detailed in Section 5.1, while light field reconstruction results are discussed in Section 5.2.
5.1. View Synthesis
| Dataset | Method | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| SinMPI (Pu et al., 2023) | 17.3295 | 0.4193 | 0.7064 | 0.4911 | 0.2794 | 0.4097 | |
| Blender | TMPI (Khan et al., 2023) | 19.9507 | 0.7160 | 0.8264 | 0.4822 | 0.1458 | 0.1471 |
| 256x256 | AdaMPI (Han et al., 2022) | 24.9252 | 0.8092 | 0.9145 | 0.6720 | 0.0680 | 0.0455 |
| PLFNet+ (Ours) | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 | |
| SinMPI (Pu et al., 2023) | 18.3441 | 0.4599 | 0.6885 | 0.3621 | 0.2251 | 0.3966 | |
| Blender | TMPI (Khan et al., 2023) | 20.0126 | 0.7071 | 0.8070 | 0.3605 | 0.1268 | 0.1640 |
| 512x512 | AdaMPI (Han et al., 2022) | 23.1416 | 0.7510 | 0.8566 | 0.4353 | 0.0751 | 0.0770 |
| PLFNet+ (Ours) | 26.3339 | 0.8334 | 0.9307 | 0.5633 | 0.2191 | 0.1947 | |
| SinMPI (Pu et al., 2023) | 14.5548 | 0.4271 | 0.6621 | 0.3706 | 0.2867 | 0.4024 | |
| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 17.1759 | 0.5454 | 0.7314 | 0.3819 | 0.2028 | 0.2415 |
| 256x256 | AdaMPI (Han et al., 2022) | 23.6155 | 0.8263 | 0.9345 | 0.7232 | 0.1055 | 0.0797 |
| PLFNet+ (Ours) | 22.8933 | 0.7608 | 0.9070 | 0.5534 | 0.2378 | 0.2560 | |
| SinMPI (Pu et al., 2023) | 15.7272 | 0.4740 | 0.6712 | 0.3095 | 0.2392 | 0.3775 | |
| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 18.1375 | 0.5444 | 0.7499 | 0.2906 | 0.1750 | 0.2758 |
| 512x512 | AdaMPI (Han et al., 2022) | 21.5931 | 0.7425 | 0.8925 | 0.5215 | 0.1156 | 0.1559 |
| PLFNet+ (Ours) | 20.2596 | 0.6229 | 0.8001 | 0.3007 | 0.2758 | 0.4226 |
| Method | 256x256 | 512x512 |
| SinMPI (Pu et al., 2023) | 0.0752 / 13 | 0.2368 / 4 |
| TMPI (Khan et al., 2023) | 0.0910 /11 | 0.2640 / 4 |
| AdaMPI (Han et al., 2022) | 0.2261 / 4 | 0.3254 / 3 |
| PLFNet+ (Ours) | 0.0074 / 135 | 0.0086 / 116 |
5.1.1. Datasets
To evaluate the performance of PLFNet+ on view synthesis, we used two datasets. The first is a custom Blender dataset, created specifically for this task, which contains images from different scenes with camera movements constrained within a 20 cm cubical volume. The second dataset is COCO (Caesar et al., 2018), which we use during training in the same manner as AdaMPI (Han et al., 2022), where depth-image-based rendering (DIBR) with hole filling is employed to generate the training data. During evaluation, we take an input image and synthesize a novel image at a specified target location. The synthesized image is then treated as input to generate an image back at the original location. This process enables us to compare the reconstructed image with the ground truth, which is the original input image, allowing for an assessment of rendering quality. Both datasets are evaluated at resolutions of and and a movement within a volume of .
5.1.2. Evaluation Metrics
To objectively evaluate the quality of the synthesized novel views, we employed multiple standard and advanced metrics. These include peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) (Wang et al., 2004), as well as multi-scale structural similarity index measure (MS-SSIM) (Wang et al., 2003), pixel- based visual information fidelity (VIFP) (Sheikh and Bovik, 2006), deep image structure and texture similarity (DISTS) (Ding et al., 2020), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018). These metrics capture not only pixel-level errors but also perceptual differences, providing a holistic view of the synthesis quality.
5.1.3. Results
Quantitative Results: The quantitative results for view synthesis across different datasets and resolutions are presented in Table 1. Our proposed PLFNet+ consistently outperforms AdaMPI (Han et al., 2022) on the Blender dataset and achieves competitive results on the COCO (Caesar et al., 2018) dataset. Although AdaMPI demonstrates slightly better performance in perceptual metrics such as DISTS and LPIPS, our method remains competitive across a range of metrics, especially on PSNR and SSIM.
Interestingly, we observe a slight performance drop in AdaMPI and PLFNet+ when increasing the resolution from to . This highlights the challenges of scaling models without increasing their capacity to handle higher-resolution inputs. In contrast to other methods that explicitly rely on depth estimation using the DPT model, our approach learns scene depth implicitly from RGB input alone, demonstrating the flexibility of our method without the need for pre-computed depth maps.
When evaluating inference time, PLFNet+ excels, particularly at higher resolutions. As shown in Table 2, our method achieves an inference time of 0.0086 seconds for a resolution of , outperforming all other methods by a significant margin. This makes PLFNet+ highly suitable for real-time applications, where rapid generation of novel views is critical, such as in dynamic video streams or interactive virtual environments.
Qualitative Results: Figure 4 provides a visual comparison of the view synthesis results for different methods on the COCO dataset. Each method’s output is accompanied by error maps, allowing for a more detailed analysis of their accuracy.
SinMPI (Pu et al., 2023) struggles with rendering artifacts, particularly in regions with complex textures and geometry, as evidenced by the error maps. These issues are likely a result of its depth estimation and scene representation. TMPI (Khan et al., 2023) also displays notable tiling artifacts due to its approach of partitioning the scene into multiple planes, leading to visible seams and discontinuities.
In contrast, both AdaMPI and our proposed PLFNet+ produce the most visually accurate results, with minimal artifacts. AdaMPI shows slightly better performance in some scenarios, particularly in handling fine details, but our method’s ability to efficiently handle reflections, occlusions, and fine textures stands out. While AdaMPI may have a slight edge in perceptual quality, our method’s superior inference speed makes it the more practical choice for real-time applications, balancing high-quality synthesis with the demands of time-sensitive environments.
5.2. Extending View Synthesis: Light Field Reconstruction
5.2.1. Datasets
For LF reconstruction, we evaluated our method using three publicly available datasets: Lytro Flowers (Srinivasan et al., 2017), Stanford Light Field Archive (Raj et al., 2016), and JPEG-Pleno (Rerabek and Ebrahimi, 2016). These datasets are widely used for benchmarking light field reconstruction algorithms. The extracted light field images had dimensions of 352x512 with a 7x7 angular resolution.
We trained PLFNet on 75% of the Lytro Flowers and Stanford datasets, with the remaining 25% reserved for testing. The JPEG-Pleno dataset, however, was exclusively used for testing to evaluate the generalization capability of our model, as it was not included in the training process.
5.2.2. Evaluation Metrics
We utilize similar evaluation metrics as described in Section 5.1.2. Since the light field reconstruction works in very narrow baseline between each sub-aperture image, and has discrete camera locations, the search space is much smaller therefore we get much better scores in these metrics.
| Dataset | Method | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| Flowers (Srinivasan et al., 2017) | NoisyLFRecon | 39.9500 | 0.9763 | 0.9932 | 0.9285 | 0.0389 | 0.0190 |
| DGLF | 35.6194 | 0.8773 | 0.9301 | 0.6021 | 0.1556 | 0.1431 | |
| DALF | 37.3006 | 0.8941 | 0.9589 | 0.7362 | 0.1033 | 0.0911 | |
| IR-V | 37.9034 | 0.9122 | 0.9645 | 0.7324 | 0.0994 | 0.0707 | |
| LFSphereNet | 41.3719 | 0.9461 | 0.9868 | 0.9060 | 0.0812 | 0.0512 | |
| PLFNet | 37.2606 | 0.9601 | 0.9914 | 0.9062 | 0.0608 | 0.0229 | |
| Stanford (Raj et al., 2016) | NoisyLFRecon | 35.4917 | 0.9572 | 0.9771 | 0.8380 | 0.0718 | 0.0395 |
| DGLF | 35.6509 | 0.9673 | 0.9897 | 0.8655 | 0.0685 | 0.0321 | |
| DALF | 38.2996 | 0.9172 | 0.9750 | 0.7829 | 0.0667 | 0.0402 | |
| IR-V | 39.4909 | 0.9440 | 0.9852 | 0.8358 | 0.0628 | 0.0390 | |
| LFSphereNet | 40.9830 | 0.9488 | 0.9797 | 0.8842 | 0.0556 | 0.0300 | |
| PLFNet | 36.8530 | 0.9601 | 0.9927 | 0.8679 | 0.0553 | 0.0182 | |
| JPEG Pleno (Rerabek and Ebrahimi, 2016) | NoisyLFRecon | 36.6275 | 0.9575 | 0.9687 | 0.8549 | 0.0718 | 0.0435 |
| DGLF | 32.3387 | 0.7207 | 0.8166 | 0.4246 | 0.1556 | 0.1431 | |
| DALF | 35.0285 | 0.8257 | 0.9291 | 0.6462 | 0.1095 | 0.0937 | |
| IR-V | 37.2085 | 0.9122 | 0.9719 | 0.7878 | 0.0894 | 0.0605 | |
| LFSphereNet | 39.2624 | 0.9429 | 0.9879 | 0.9069 | 0.0737 | 0.0425 | |
| PLFNet | 33.8982 | 0.9106 | 0.9838 | 0.8683 | 0.0590 | 0.0261 |
| Method | Purpose | Inference Time (s) |
| NoisyLFRecon | LF reconstruction | 1.1830 |
| DGLF | LF reconstruction | 1.5009 |
| DALF | LF reconstruction | 0.0583 |
| IR-V | LF reconstruction | 0.1429 |
| LFSphereNet | LF reconstruction | 0.0008 |
| PLFNet | Novel view synthesis | 0.0980 |
| NoisyLFRecon | DGLF | DALF | IR-V | LFSphereNet | PLFNet |
| (Zhou et al., 2021) | (Zhou et al., 2020) | (Cun et al., 2019) | (Han and Xiang, 2022) | (Gond et al., 2023) | (Ours) |
5.2.3. Results
Quantitative Results: The quantitative results of our light field reconstruction experiments are presented in Table 3. Our proposed PLFNet demonstrates competitive performance compared to state-of-the-art methods, including NoisyLFRecon (Zhou et al., 2021), DGLF (Zhou et al., 2020), DALF (Cun et al., 2019), IR-V (Han and Xiang, 2022), and LFSphereNet (Gond et al., 2023). All models were trained from scratch using identical datasets for a fair comparison. Publicly available implementations were used for the methods by Zhou et al. (2021, 2020) and Han and Xiang (2022), while DALF (Cun et al., 2019) was re-implemented in PyTorch according to the original paper.
Although DGLF and LFSphereNet generate the full light field in a single pass, both DALF and our PLFNet reconstruct each sub-aperture image iteratively. For the Stanford dataset, which offers fewer training samples compared to Lytro Flowers, LFSphereNet delivered the best results across most metrics, with PLFNet showing competitive performance, particularly excelling in SSIM and MS-SSIM. In terms of computational efficiency, PLFNet demonstrated significant improvements as shown in Table 4. Our PLFNet achieved an inference time of 0.0980 seconds for whole LF reconstruction which included forward passes. However, per view inference on average is around 0.006 seconds highlighting its potential for real-time applications.
Qualitative Results: The qualitative evaluation of light field reconstruction is illustrated in Figure 5. We compare the visual performance of our proposed PLFNet against state-of-the-art methods for light field views, focusing on the top right corner sub-aperture view. The first two rows depict reconstruction results for the JPEG-Pleno (Rerabek and Ebrahimi, 2016) dataset. In these challenging scenarios, PLFNet demonstrates superior reconstruction quality with fewer visible artifacts, as evident in the error maps. Compared to LFSphereNet (Gond et al., 2023), which also shows strong performance, PLFNet exhibits more accurate details and reduced error magnitudes, highlighting its robustness and generalization capabilities. The subsequent rows display results for the Stanford Light Field Archive (Raj et al., 2016) dataset. Here, our method consistently outperforms NoisyLFRecon (Zhou et al., 2021), which utilizes nine input images for reconstruction. Despite this advantage, NoisyLFRecon produces noticeable artifacts and blurring in regions with fine structures. In contrast, PLFNet, which relies on only a single input image, achieves sharper reconstructions with more faithful reproduction of scene geometry and textures. This indicates that PLFNet can effectively capture and synthesize complex light field information even with minimal input data.
Overall, the qualitative results corroborate the quantitative findings, showcasing ability of PLFNet to achieve high-quality reconstructions with low error rates. This performance shows potential for practical applications where high-fidelity light field reconstruction from sparse data is required.
6. Ablation Study
In this section, we present a two-part ablation study to assess the effectiveness of our proposed method, focusing on positional understanding and the impact of different positional embedding techniques. The first part evaluates the performance of PLFNet+ under varying sparse input configurations, demonstrating how the number and location of input images influence view synthesis. The second part investigates the contribution of different positional encoding strategies to the overall reconstruction quality, highlighting the role of each component in our final model.
6.1. Sparse Input Mode
The first part of the ablation study explores how the performance of PLFNet+ is affected by varying the number and position of input images for view synthesis. We evaluate three configurations: (i) a single central image as input, (ii) a single image from a random location within the cubical volume, and (iii) two images from random locations. These setups allow us to understand how well the model generalizes to sparse input scenarios, particularly in cases where inputs are not centrally aligned.
Table 5 presents the quantitative results of this study. As expected, the single central view input yields the best performance across all metrics, with a PSNR of 28.8023 and SSIM of 0.9009, indicating that the model performs best when provided with a centrally located reference image. The random input configurations, whether with one or two images, exhibit slightly reduced performance, suggesting that while the model can generalize to inputs from random locations, the quality of the synthesized view degrades with less structured input data. Nevertheless, even with these random inputs, PLFNet+ achieves acceptable performance, demonstrating its robustness in handling diverse input scenarios.
| Number of Input Images | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| 2 Images - Random Location | 23.0571 | 0.7444 | 0.8462 | 0.4297 | 0.2782 | 0.2510 |
| 1 Image - Random Location | 22.8429 | 0.7243 | 0.8386 | 0.3524 | 0.3373 | 0.3703 |
| 1 Image - Central View | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 |
6.2. Positional Embedding
In the second part of the ablation study, we analyze the impact of different positional embedding strategies on view synthesis. We compare four configurations: (i) MLP only, (ii) normalization only, (iii) normalization with positional encoding, (iv) normalization with MLP, and (v) the full setup with normalization, positional encoding, and MLP. This analysis aims to determine the contribution of each component in enhancing the model’s understanding of positional information.
The results are summarized in Table 6. The combination of normalization, positional encoding, and MLP consistently outperforms the other configurations, achieving a PSNR of 28.8023 and SSIM of 0.9003. This demonstrates the effectiveness of our final proposed method in capturing positional variations and synthesizing accurate novel views. By contrast, using only normalization or MLP leads to a noticeable drop in performance specially in high frequency regions, underscoring the importance of combining these techniques to achieve optimal results.
The setup with normalization and MLP alone performs better than using MLP in isolation but still falls short of the complete configuration. These findings highlight the critical role of positional encoding and MLP in providing a richer representation of the 3D positional information, which is essential for accurately reconstructing complex novel views.
| Positional Embedding | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| MLP | 23.2578 | 0.7633 | 0.8686 | 0.4063 | 0.2566 | 0.2726 |
| Normalization | 25.2855 | 0.8376 | 0.9256 | 0.5309 | 0.2132 | 0.1680 |
| Norm+Positional Encoding | 24.1945 | 0.7941 | 0.8962 | 0.4502 | 0.2276 | 0.1946 |
| Norm+MLP | 26.4475 | 0.8508 | 0.9421 | 0.6247 | 0.2119 | 0.1647 |
| Norm+PosEnc+MLP | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 |
7. Conclusion
In this paper we presented a novel, lightweight, position-aware network for real-time view synthesis and light field reconstruction, which leverages a position-aware embedding combined with a rendering network to achieve both efficiency and high visual quality. Through the development of PLFNet and its extension, PLFNet+, our approach demonstrates superior synthesis quality, particularly in handling complex translational movements without relying on computationally intensive operations like warping.
Extensive evaluations reveal that PLFNet+ achieves a strong balance between visual fidelity and inference speed, positioning it as a feasible solution for applications requiring real-time performance, such as telepresence and augmented reality. Additionally, our network architecture shows robustness in various scenarios, from narrow to wide baseline settings, making it adaptable to dynamic and interactive environments.
The promising results from ablation studies underscore the effectiveness of our positional embedding design, particularly the integration of normalization, positional encoding, and MLP components. Overall, our work provides a significant step towards more accessible, high-performance view synthesis frameworks, paving the way for broader application in consumer and industrial devices.
Acknowledgements.
The work was supported by the European Joint Doctoral Programme on Plenoptic Imaging (PLENOPTIMA) through the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 956770. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. We thank the High Performance Computing Center North (HPC2N) at Umeå University for providing computational resources and valuable support during test and performance runs.References
- (1)
- Bae et al. (2021) Kyuho Bae, Andre Ivan, Hajime Nagahara, and In Kyu Park. 2021. 5d light field synthesis from a monocular video. In 2020 25th International Conference on Pattern Recognition (ICPR). IEEE, 7157–7164.
- Barron et al. (2021) Jonathan T Barron, Ben Mildenhall, Matthew Tancik, Peter Hedman, Ricardo Martin-Brualla, and Pratul P Srinivasan. 2021. Mip-nerf: A multiscale representation for anti-aliasing neural radiance fields. In Proceedings of the IEEE/CVF international conference on computer vision. 5855–5864.
- Caesar et al. (2018) Holger Caesar, Jasper Uijlings, and Vittorio Ferrari. 2018. Coco-stuff: Thing and stuff classes in context. In Proceedings of the IEEE conference on computer vision and pattern recognition. 1209–1218.
- Chandramouli et al. (2020) Paramanand Chandramouli, Kanchana Vaishnavi Gandikota, Andreas Goerlitz, Andreas Kolb, and Michael Moeller. 2020. A generative model for generic light field reconstruction. IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 4 (2020), 1712–1724.
- Chen et al. (2020) Bin Chen, Lingyan Ruan, and Miu-Ling Lam. 2020. LFGAN: 4D Light Field Synthesis from a Single RGB Image. ACM Trans. Multimedia Comput. Commun. Appl. 16 (2 2020). Issue 1. https://doi.org/10.1145/3366371
- Cun et al. (2019) Xiaodong Cun, Feng Xu, Chi-Man Pun, and Hao Gao. 2019. Depth-Assisted Full Resolution Network for Single Image-Based View Synthesis. IEEE Computer Graphics and Applications 39 (2019), 52–64. Issue 2. https://doi.org/10.1109/MCG.2018.2884188
- Deitke et al. (2023) Matt Deitke, Dustin Schwenk, Jordi Salvador, Luca Weihs, Oscar Michel, Eli VanderBilt, Ludwig Schmidt, Kiana Ehsani, Aniruddha Kembhavi, and Ali Farhadi. 2023. Objaverse: A universe of annotated 3d objects. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 13142–13153.
- Dima and Sjöström (2021) Elijs Dima and Mårten Sjöström. 2021. Camera and Lidar-Based View Generation for Augmented Remote Operation in Mining Applications. IEEE Access 9 (2021), 82199–82212.
- Ding et al. (2020) Keyan Ding, Kede Ma, Shiqi Wang, and Eero P Simoncelli. 2020. Image quality assessment: Unifying structure and texture similarity. IEEE transactions on pattern analysis and machine intelligence 44, 5 (2020), 2567–2581.
- Dosovitskiy (2020) Alexey Dosovitskiy. 2020. An image is worth 16x16 words: Transformers for image recognition at scale. arXiv preprint arXiv:2010.11929 (2020).
- Flynn et al. (2019) John Flynn, Michael Broxton, Paul Debevec, Matthew DuVall, Graham Fyffe, Ryan Overbeck, Noah Snavely, and Richard Tucker. 2019. DeepView: View Synthesis With Learned Gradient Descent. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR).
- Fridovich-Keil et al. (2022) Sara Fridovich-Keil, Alex Yu, Matthew Tancik, Qinhong Chen, Benjamin Recht, and Angjoo Kanazawa. 2022. Plenoxels: Radiance fields without neural networks. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 5501–5510.
- Gond et al. (2023) Manu Gond, Emin Zerman, Sebastian Knorr, and Mårten Sjöström. 2023. LFSphereNet: Real Time Spherical Light Field Reconstruction from a Single Omnidirectional Image. In Proceedings of the 20th ACM SIGGRAPH European Conference on Visual Media Production. 1–10.
- Han and Xiang (2022) Kang Han and Wei Xiang. 2022. Inference-Reconstruction Variational Autoencoder for Light Field Image Reconstruction. IEEE Transactions on Image Processing 31 (2022), 5629–5644. https://doi.org/10.1109/TIP.2022.3197976
- Han et al. (2022) Yuxuan Han, Ruicheng Wang, and Jiaolong Yang. 2022. Single-view view synthesis in the wild with learned adaptive multiplane images. In ACM SIGGRAPH 2022 Conference Proceedings. 1–8.
- He et al. (2016) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016. Deep residual learning for image recognition. In Proceedings of the IEEE conference on computer vision and pattern recognition. 770–778.
- Isaksen et al. (2000) Aaron Isaksen, Leonard McMillan, and Steven J Gortler. 2000. Dynamically reparameterized light fields. In Proceedings of the 27th annual conference on Computer graphics and interactive techniques. 297–306.
- Ivan et al. (2019) Andre Ivan, In Kyu Park, et al. 2019. Synthesizing a 4D spatio-angular consistent light field from a single image. arXiv preprint arXiv:1903.12364 (2019).
- Jang and Agapito (2024) Wonbong Jang and Lourdes Agapito. 2024. NViST: In the Wild New View Synthesis from a Single Image with Transformers. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 10181–10193.
- Jiang et al. (2021) Liming Jiang, Bo Dai, Wayne Wu, and Chen Change Loy. 2021. Focal frequency loss for image reconstruction and synthesis. In Proceedings of the IEEE/CVF international conference on computer vision. 13919–13929.
- Kerbl et al. (2023) Bernhard Kerbl, Georgios Kopanas, Thomas Leimkühler, and George Drettakis. 2023. 3D Gaussian Splatting for Real-Time Radiance Field Rendering. ACM Trans. Graph. 42, 4 (2023), 139–1.
- Khan et al. (2023) Numair Khan, Lei Xiao, and Douglas Lanman. 2023. Tiled multiplane images for practical 3d photography. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 10454–10464.
- Koh et al. (2023) Jing Yu Koh, Harsh Agrawal, Dhruv Batra, Richard Tucker, Austin Waters, Honglak Lee, Yinfei Yang, Jason Baldridge, and Peter Anderson. 2023. Simple and effective synthesis of indoor 3d scenes. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 37. 1169–1178.
- Levoy and Hanrahan (1996) Marc Levoy and Pat Hanrahan. 1996. Light field rendering. In Proceedings of the 23rd annual conference on Computer graphics and interactive techniques. 31–42.
- Li and Kalantari (2020) Qinbo Li and Nima Khademi Kalantari. 2020. Synthesizing Light Field from a Single Image with Variable MPI and Two Network Fusion. ACM Trans. Graph. 39 (11 2020). Issue 6. https://doi.org/10.1145/3414685.3417785
- Lin et al. (2023) Kai-En Lin, Yen-Chen Lin, Wei-Sheng Lai, Tsung-Yi Lin, Yi-Chang Shih, and Ravi Ramamoorthi. 2023. Vision transformer for nerf-based view synthesis from a single input image. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 806–815.
- Liu et al. (2024) Minghua Liu, Chao Xu, Haian Jin, Linghao Chen, Mukund Varma T, Zexiang Xu, and Hao Su. 2024. One-2-3-45: Any single image to 3d mesh in 45 seconds without per-shape optimization. Advances in Neural Information Processing Systems 36 (2024).
- Liu et al. (2023) Ruoshi Liu, Rundi Wu, Basile Van Hoorick, Pavel Tokmakov, Sergey Zakharov, and Carl Vondrick. 2023. Zero-1-to-3: Zero-shot one image to 3d object. In Proceedings of the IEEE/CVF international conference on computer vision. 9298–9309.
- Mildenhall et al. (2021) Ben Mildenhall, Pratul P. Srinivasan, Matthew Tancik, Jonathan T. Barron, Ravi Ramamoorthi, and Ren Ng. 2021. NeRF: Representing Scenes as Neural Radiance Fields for View Synthesis. Commun. ACM 65, 1 (dec 2021), 99–106. https://doi.org/10.1145/3503250
- Müller et al. (2022) Thomas Müller, Alex Evans, Christoph Schied, and Alexander Keller. 2022. Instant neural graphics primitives with a multiresolution hash encoding. ACM transactions on graphics (TOG) 41, 4 (2022), 1–15.
- Overbeck et al. (2018) Ryan S Overbeck, Daniel Erickson, Daniel Evangelakos, Matt Pharr, and Paul Debevec. 2018. A system for acquiring, processing, and rendering panoramic light field stills for virtual reality. ACM Transactions on Graphics (TOG) 37, 6 (2018), 1–15.
- Pu et al. (2023) Guo Pu, Peng-Shuai Wang, and Zhouhui Lian. 2023. SinMPI: Novel View Synthesis from a Single Image with Expanded Multiplane Images. In SIGGRAPH Asia 2023 Conference Papers. 1–10.
- Raj et al. (2016) Abhilash Sunder Raj, Michael Lowney, Raj Shah, and Gordon Wetzstein. 2016. Stanford lytro light field archive.
- Rerabek and Ebrahimi (2016) Martin Rerabek and Touradj Ebrahimi. 2016. New light field image dataset. In 8th International Conference on Quality of Multimedia Experience (QoMEX).
- Seitz and Dyer (1999) Steven M Seitz and Charles R Dyer. 1999. Photorealistic scene reconstruction by voxel coloring. International journal of computer vision 35 (1999), 151–173.
- Sheikh and Bovik (2006) Hamid R Sheikh and Alan C Bovik. 2006. Image information and visual quality. IEEE Transactions on image processing 15, 2 (2006), 430–444.
- Srinivasan et al. (2017) Pratul P Srinivasan, Tongzhou Wang, Ashwin Sreelal, Ravi Ramamoorthi, and Ren Ng. 2017. Learning to Synthesize a 4D RGBD Light Field From a Single Image. Proceedings of the IEEE International Conference on Computer Vision (ICCV). https://openaccess.thecvf.com/content_iccv_2017/html/Srinivasan_Learning_to_Synthesize_ICCV_2017_paper.html
- Tucker and Snavely (2020) Richard Tucker and Noah Snavely. 2020. Single-view view synthesis with multiplane images. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 551–560.
- Wang et al. (2004) Zhou Wang, A.C. Bovik, H.R. Sheikh, and E.P. Simoncelli. 2004. Image quality assessment: from error visibility to structural similarity. IEEE Transactions on Image Processing 13, 4 (2004), 600–612. https://doi.org/10.1109/TIP.2003.819861
- Wang et al. (2003) Zhou Wang, Eero P Simoncelli, and Alan C Bovik. 2003. Multiscale structural similarity for image quality assessment. In The Thrity-Seventh Asilomar Conference on Signals, Systems & Computers, 2003, Vol. 2. Ieee, 1398–1402.
- Watson et al. (2022) Daniel Watson, William Chan, Ricardo Martin-Brualla, Jonathan Ho, Andrea Tagliasacchi, and Mohammad Norouzi. 2022. Novel view synthesis with diffusion models. arXiv preprint arXiv:2210.04628 (2022).
- Yu et al. (2021) Alex Yu, Vickie Ye, Matthew Tancik, and Angjoo Kanazawa. 2021. pixelnerf: Neural radiance fields from one or few images. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 4578–4587.
- Zerman et al. (2024) Emin Zerman, Manu Gond, Soheib Takhtardeshir, Roger Olsson, and Mårten Sjöström. 2024. A Spherical Light Field Database for Immersive Telecommunication and Telepresence Applications. In 2024 16th International Conference on Quality of Multimedia Experience (QoMEX). 200–206. https://doi.org/10.1109/QoMEX61742.2024.10598264
- Zhang et al. (2018) Richard Zhang, Phillip Isola, Alexei A Efros, Eli Shechtman, and Oliver Wang. 2018. The Unreasonable Effectiveness of Deep Features as a Perceptual Metric. In CVPR.
- Zhao et al. (2016) Hang Zhao, Orazio Gallo, Iuri Frosio, and Jan Kautz. 2016. Loss functions for image restoration with neural networks. IEEE Transactions on computational imaging 3, 1 (2016), 47–57.
- Zhou et al. (2018) Tinghui Zhou, Richard Tucker, John Flynn, Graham Fyffe, and Noah Snavely. 2018. Stereo Magnification: Learning View Synthesis Using Multiplane Images. ACM Trans. Graph. 37, 4, Article 65 (jul 2018), 12 pages. https://doi.org/10.1145/3197517.3201323
- Zhou et al. (2020) Wenhui Zhou, Gaomin Liu, Jiangwei Shi, Hua Zhang, and Guojun Dai. 2020. Depth-guided view synthesis for light field reconstruction from a single image. Image and Vision Computing 95 (2020), 103874. https://doi.org/10.1016/j.imavis.2020.103874
- Zhou et al. (2021) Wenhui Zhou, Jiangwei Shi, Yongjie Hong, Lili Lin, and Ercan Engin Kuruoglu. 2021. Robust dense light field reconstruction from sparse noisy sampling. Signal Processing 186 (9 2021). https://doi.org/10.1016/j.sigpro.2021.108121

### v2

@@
 Abstract.
-Recent advancements in view synthesis have significantly enhanced immersive experiences across various computer graphics and multimedia applications, including telepresence, and entertainment.
+Recent advancements in view synthesis have significantly enhanced immersive experiences across various computer graphics and multimedia applications, including telepresence and entertainment.
 By enabling the generation of new perspectives from a single input view, view synthesis allows users to better perceive and interact with their environment.
@@
 In this paper, we present a lightweight, position-aware network designed for real-time view synthesis from a single input image and a target camera pose.
-The proposed framework consists of a Position Aware Embedding, modeled with a multi-layer perceptron, which efficiently maps positional information from the target pose to generate high dimensional feature maps.
-These feature maps, along with the input image, are fed into a Rendering Network that merges features from dual encoder branches to resolve both high level semantics and low level details, producing a realistic new view of the scene.
+The proposed framework consists of a Position Aware Embedding, which efficiently maps positional information from the target pose to generate high dimensional feature maps.
+These feature maps, along with the input image, are fed into a Rendering Network that merges features from dual encoder branches to resolve both high and low level details, producing a realistic new view of the scene.
 Experimental results demonstrate that our method achieves superior efficiency and visual quality compared to existing approaches, particularly in handling complex translational movements without explicit geometric operations like warping.
-This work marks a step toward enabling real-time view synthesis from a single image for live and interactive applications.
+This work marks a step toward enabling real-time live and interactive telepresence applications.
 1. Introduction
-Telepresence systems (Dima and Sjöström, 2021) enable users to interact with remote environments in real time, offering a wide range of applications, particularly in sectors such as heavy machinery, where remote operation from a safe distance is crucial.
+Telepresence systems (Dima and Sjöström, 2021; Chandan et al., 2021; Sakashita et al., 2023), allow users to interact with remote environments in real-time, offering a wide range of applications, particularly in sectors such as heavy machinery, where remote operation from a safe distance is crucial.
 These systems have the potential to facilitate seamless natural interaction with remote environments.
-Current telepresence technologies typically rely on head-mounted displays (HMDs), high-speed networks, and consumer or industrial-grade cameras to deliver live video feeds of the environment.
-While 2D video feeds provide some pictorial information, the lack of depth perception significantly limits situational awareness, which is vital for effective remote operation.
-The challenge to have an understanding of a remote site’s structure may be address by enabling motion parallax for the user.
-I.e., the user moves the head forth and back to get a notion of the scene depth.
-This requires access to views from different perspectives, where camera positions and poses need to be continuous for a smooth transition and a high quality experience.
-Therefore, it is imperative to render continuous perspective views accurately and in real time for applications such as augmented reality (AR), virtual reality (VR), and telepresence systems.
-The synthesis of views from perspectives continuous in camera position and pose requires sufficient information of the scene.
-Early methods employed view interpolations and depth-image-based rendering, but both these require multiple capture devices resulting in calibration issues and high communication capacity requirements.
+Current real-time telepresence technologies (Chandan et al., 2021; Sakashita et al., 2023) typically rely on head-mounted displays (HMDs), high-speed networks, and consumer or industrial-grade cameras to deliver live video feeds of the environment.
+While live 2D video feeds provide visual information, they fall short in offering depth perception (Chandan et al., 2021).
+The lack of depth perception significantly limits situational awareness (Dima et al., 2020), which is vital for effective remote operation.
+The lack of depth perception that a 2D video feed provides can be leveraged by binocular parallax and motion parallax as depth cues.
+Recently, SpinVR (Konrad et al., 2017) utilized a stereoscopic virtual reality setup to provide binocular parallax to the users.
+However, this still lacks motion parallax.
+When a user moves their viewpoint in space, the relative motion of objects within the scene provides the brain with cues about the depth.
+Thus, rendering new views at different positions can simulate motion parallax, improving depth perception and enhancing situational awareness in remote operations.
+For effective depth perception and smooth interaction in real-time, synthesizing new views, while maintaining low latency, becomes essential as it enables both binocular and motion parallaxes.
+Synthesizing views from different perspectives requires sufficient information of the scene.
+While 360-degree cameras are used in some telepresence system (Chandan et al., 2021), synthesizing new views from these omnidirectional images (ODIs) as done in works like PanoSynthVR (Waidhofer et al., 2022) introduces additional complexities.
+The distortion inherent in ODIs and the required high bandwidth for transferring such data make these methods less practical for real-time use.
+Instead, focusing on planar images captured from the viewpoint of the user offers a more efficient solution.
+We consider a telepresence application as the primary use case for our view synthesis method, illustrated in Figure 1.
+In this setting, an omnidirectional image (ODI) is captured from remote scene.
+A viewport image is extracted from the ODI based on the current gaze direction of the user, while the relative head movement (i.e., translation) is used to define a target position.
+The extracted image, target position, and target rotation are then passed to our network, which synthesizes a new view at the intended offset.
+The resulting image is rendered as the updated viewport, enabling motion parallax without requiring a full 3D reconstruction or stitching it back into the ODI.
+This patch based rendering pipeline eliminates the distortions which are present in ODIs, reduces transmission bandwidth, and allows for efficient local view synthesis.
+By synthesizing views directly from these extracted planar patches, we avoid the distortions inherent in ODIs and reduce the bandwidth needed for transmitting high resolution spherical imagery.
+This patch based synthesis pipeline supports real-time performance and enables localized motion parallax around the viewpoint of the user.
+However, such ODI-based workflows inherently embed orientation within the extracted view direction, requiring only translational offsets for rendering.
+To extend support to general planar RGB inputs such as monocular cameras not embedded in a spherical system, our method also incorporates full 6DoF inputs, including yaw, pitch, and roll.
+This makes our framework broadly applicable beyond the illustrated telepresence scenario.
+Early view synthesis methods (Ahn and Kim, 2013; Yao et al., 2018; Zhu et al., 2019) employed view interpolations (Chen and Williams, 1993) and depth-image-based rendering (Fehn, 2004) , but these suffer from interpolation artifacts, and pixel misalignment due to the use of noisy depth maps.
+Recent methods like Quark (Flynn et al., 2024) and SIMPLI (Solovev et al., 2023) employ a dynamic layered depth map (LDM) of the scene estimated through a deep neural network to perform view synthesis.
+However, these methods still require multi-view input, resulting in requirement of multiple capture devices and high communication bandwidth.
 Therefore, generating new views from a single input image and a specified camera pose has emerged as a key task in rendering.
-A variety of approaches, such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and more recent transformer-based architectures like NViST (Jang and Agapito, 2024), have been proposed to tackle this problem.
-These methods leverage both synthetic and real-world datasets to produce high-quality novel views.
-However, most state-of-the-art methods either require multi-view inputs or are computationally intensive, making them unsuitable for real-time applications where low latency is critical.
-In this work, we build upon the growing body of research on single-image-based view synthesis.
-We introduce a learning based position embedding method combined with a rendering network that efficiently synthesizes high-quality views of arbitrary camera location and pose using fewer computational resources.
-Our method maintains high fidelity even when handling large translational shifts, making it ideal for real-time applications and practical deployment on consumer devices.
-To train and test our method we also create our own synthetic dataset using Blender.
-Unlike other methods, such as NViST, which employs transformer architectures, or SE3D (Koh et al., 2023), which utilizes large-scale generative-adversarial networks (GANs) to address occlusions and complex scenes, our network is designed to be lightweight and computationally efficient.
-Despite its smaller footprint, it excels in handling scenarios where translational movements dominate, enabling real-time inference.
+A variety of approaches, diffusion based 3DiM (Watson et al., 2022) and transformer-based architectures like NViST (Jang and Agapito, 2024), are proposed to tackle this problem.
+These methods leverage large scale datasets to be trained and produce high-quality novel views.
+While these approaches achieve remarkable fidelity, they are either unsuitable for real-time deployment or assume specific scene representations (e.g., dense layers or precomputed geometry).
+Our work builds on this trajectory by focusing on generalizable, feed forward synthesis that supports both real-time constraints and 6DoF pose inputs.
+Real-time rendering is critical for telepresence and AR applications, but most state-of-the-art methods either depend on heavy iterative models (e.g., diffusion (Watson et al., 2022; Deitke et al., 2023)) or rely on attention-heavy transformer blocks (Jang and Agapito, 2024), both of which impede latency-sensitive deployments.
+In this work, we build upon the growing body of research to address the challenges of computation efficiency and low latency for view synthesis.
+We propose a method designed to generate high-quality novel views from a single input image, named PVSNet, targeting real-time performance.
+PVSNet utilizes a position-aware rendering approach that efficiently handles camera movements in the scene, enabling smooth and realistic motion parallax for depth perception without the computational burden of traditional methods.
 In summary, our main contributions are as follows:
-A lightweight network with a positional embedding module that can generate novel views within a given baseline without requiring explicit operations.
-A positional encoding scheme similar to NeRF, which is combined with an multi-layer perceptron (MLP), to create a more generalizable framework capable of handling different scenes.
-Our method achieves real-time performance, with inference times of approximately 8.6 milliseconds, corresponding to over 100 frames per second (FPS) for images at a resolution of 512x512 pixels.
-The remainder of this paper is organized as follows: Section 2 reviews state-of-the-art methods aimed at view synthesis.
-Section 3, and 4 describes the proposed approach for small baseline and large baseline rendering.
+A neural network for real-time view synthesis from a single image and a target pose, supporting both translation only and full 6DoF inputs, making it suitable for telepresence applications.
+A compact and flexible positional embedding module similar to the one used in NeRF (Mildenhall et al., 2021), which is combined with a multi-layer perceptron (MLP), to create a more generalizable framework capable of handling different scenes.
+An in-depth evaluation of the proposed method, demonstrating real-time performance with inference times of over 50 FPS at 512512 resolution, significantly outperforming existing single view methods in speed and image quality.
+The remainder of this paper is organized as follows: Section 2 reviews the state-of-the-art methods aimed at view synthesis.
+Section 3 describes the proposed approach for view synthesis.
 Section 4 provides a comparative analysis of the results, including quality and speed evaluations.
+Section 5 presents a detailed ablation study to analyze different components of the proposed method.
 Finally, conclusions are summarized in Section 6.
 2. Background
-View synthesis has recently become a pivotal research area in immersive telepresence, aiming to generate unseen perspectives of a remote scene from a limited set of input images.
+View synthesis has recently become a pivotal research area in photorealistic immersive telepresence applications, aiming to generate unseen perspectives of a remote scene from a limited set of input images.
 Traditional approaches, such as structure-from-motion (SfM) and multi-view stereo (MVS), rely on geometric consistency between input views to reconstruct novel viewpoints (Seitz and Dyer, 1999).
 While these methods have shown effectiveness in controlled environments, their performance deteriorates in complex real-world scenarios and often necessitates multiple input views to achieve high-quality results.
-In light of these limitations, more recent methods focus on synthesizing novel views from a sparse or single image.
-This section outlines these advancements and their respective contributions to the field of view synthesis.
-2.1. Scene Regression Approaches
-Recent advancements in view synthesis have centered around methods that aim to represent a scene from sparse image inputs, particularly those leveraging neural implicit representations such as NeRF (Mildenhall et al., 2021).
+In light of these limitations, more recent methods focus on synthesizing novel views from sparse or single input images.
+This section outlines these advancements and their respective contributions to the field of view synthesis based on the number of input images required to perform view synthesis.
+2.1. Multi-View Input
+Multi-view input based methods rely on several images or views of a scene to generate novel viewpoints.
+Recent advancements in multi-view input based methods have centered around leveraging neural implicit representations such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021).
 A NeRF encodes a scene as a continuous volumetric function, allowing novel view generation through volumetric rendering.
@@
 While these improvements accelerate rendering, they retain significant reliance on MLPs, which remain computationally expensive.
-To address these concerns, methods like Plenoxels (Fridovich-Keil et al., 2022) have emerged, leveraging spherical harmonics to represent a scene without the need for deep neural networks.
-Though Plenoxels achieves faster optimization than many NeRF-based methods, it still falls short in terms of memory efficiency.
-A more recent innovation, Gaussian Splatting (Kerbl et al., 2023), improves memory and rendering efficiency by employing high-dimensional Gaussian kernels to represent the scene.
-However, its optimal performance still depends on the availability of multiple input views.
-2.2. Light Field Reconstruction and Rendering
-Light field (LF) representation encodes both the spatial and angular information of a scene, often described by an HxWxNxN structure, where HxW refers to the spatial resolution and NxN denotes the angular resolution (Levoy and Hanrahan, 1996).
+Plenoxels (Fridovich-Keil et al., 2022), which leverage spherical harmonics to represent a scene without the need for deep neural networks, have shown promising results when considering their fast optimization time.
+Though Plenoxels achieve faster optimization than many NeRF-based methods, it still falls short in terms of memory efficiency.
+A more recent innovation, Gaussian Splatting (Kerbl et al., 2023), improves rendering efficiency by employing high-dimensional Gaussian kernels to represent the scene.
+However, its optimal performance still depends on the availability of multiple input views, and the scenes with high frequency details require more kernels, increasing overall memory requirement.
+Light Field Rendering emerges as an alternative of scene regression methods like NeRF (Mildenhall et al., 2021), Plenoxels (Fridovich-Keil et al., 2022), and Gaussian Splatting (Kerbl et al., 2023).
+Light field (LF) representation encodes both the spatial and angular information of a scene, often described by an HxWxNxM structure, where HxW refers to the spatial resolution and NxM denotes the angular resolution (Levoy and Hanrahan, 1996).
 This representation facilitates tasks like viewpoint manipulation (Zhou et al., 2020).
-LF reconstruction methods aim to generate a dense NxN light field from a single input view or a few input views, eliminating the need for complex multi-camera setups.
-A common LF reconstruction approach utilizes convolutional neural networks (CNNs) to estimate scene depth (Srinivasan et al., 2017), followed by warping operators to generate target views.
-To address warping artifacts, subsequent refinement networks have been developed (Zhou et al., 2020, 2021).
-Other methods exploit depth estimation or optical flow (Cun et al., 2019), appearance flow (Ivan et al., 2019), or a combination of both (Bae et al., 2021) to synthesize the light field.
-In more recent developments, GANs (Chen et al., 2020; Chandramouli et al., 2020) and auto-encoders (Han and Xiang, 2022) have demonstrated effectiveness in reconstructing dense light fields from a single image input.
-Additionally, LFSphereNet (Gond et al., 2023) extends LF reconstruction to spherical light fields using a single omnidirectional image.
-Furthermore, to generate the novel views from the LFs, the algorithms like dynamically reparameterized LFs (Isaksen et al., 2000) and per-view disk-based blending (Overbeck et al., 2018) have to be utilized which have higher memory footprint, making the whole process unsuitable for dynamic environments and low memory systems.
-Our proposed method, PLFNet (described in Sec 3), contributes to LF reconstruction and rendering by directly synthesizing views within a local LF volume, thus eliminating the need of additional rendering algorithms.
-2.3. Learning-Based Approaches
-Learning-based methods seek to overcome the limitations of both scene regression and light field rendering approaches, which include constraints such as low baselines, sparse image input requirements, and slow rendering times.
-Several models, including PixelNeRF (Yu et al., 2021) and VisionNeRF (Lin et al., 2023), have been developed to generalize NeRF to novel scenes by conditioning the network on image features.
-Despite their promise, these methods often face challenges such as scale ambiguity and difficulty in handling large translational movements.
-To address these challenges, the diffusion model based approach in 3DiM (Watson et al., 2022) employs pose guided diffusion to improve performance, particularly in scenarios with large translational movement.
-Fine-tuning diffusion models on large-scale synthetic datasets (Deitke et al., 2023) has also shown potential (Liu et al., 2023, 2024), although these methods are hindered by slow inference times.
-An alternative strategy involves using multiplane image (MPI) representations, which approximate a scene using a series of parallel planes at fixed depths, each containing an image and alpha map.
-This representation allows novel view generation based on a fixed (Zhou et al., 2018; Flynn et al., 2019; Tucker and Snavely, 2020; Li and Kalantari, 2020) or variable (Li and Kalantari, 2020) number of planes.
+LF reconstruction methods aim to generate a dense NxM light field from few input views, eliminating the need for complex multi-camera setups.
+Common LF reconstruction methods (Zhou et al., 2020, 2021) utilize convolutional neural networks (CNNs) to estimate scene depth followed by a warping operator to reconstruct the full LF.
+To generate the novel views from the reconstructed LFs, algorithms like dynamically reparameterized LFs (Isaksen et al., 2000) and per-view disk-based blending (Overbeck et al., 2018) have to be employed.
+However, these methods often have a higher memory footprint, making the process unsuitable for dynamic environments or systems with limited memory.
+Despite the effectiveness of LF-based methods, they also face challenges when working in live applications, where real-time performance and memory efficiency are critical.
+Recent works have introduced more efficient solutions to overcome these issues, such as SIMPLI (Solovev et al., 2023), GNPR (Suhail et al., 2022), and Quark (Flynn et al., 2024).
+SIMPLI utilizes a self-improving multiplane-to-layer image approach, enhancing view synthesis by leveraging a deep refinement network for scene representation.
+However, it still faces challenges in achieving fast inference speeds while maintaining high-quality results.
+GNPR, on the other hand, uses a generalizable neural framework that extracts patches from neighboring images along the epipolar lines to generate novel views.
+This method has demonstrated impressive quality, but like SIMPLI, it struggles with high computational demands and inefficiencies during inference.
+In contrast, Quark introduces a novel layered depth map (LDM) approach that combines efficiency and quality by performing a multi-scale iterative render-and-refine process.
+This method achieves real-time, high-resolution rendering, surpassing SIMPLI and GNPR in terms of both speed and quality.
+Despite its impressive performance, Quark remains limited to multi-view inputs, restricting its applicability in scenarios where only a single input image is available.
+Our proposed method, PVSNet described in Section 3, goes a step further by offering a more efficient view synthesis approach, utilizing only a single input image.
+By removing the dependency on multiple views, PVSNet not only reduces the computational overhead, but also extends the applicability of view synthesis methods to real-time environments.
+In addition, the method ensures that high-quality novel views are generated while maintaining efficiency, making it suitable for resource-constrained applications.
+Unlike methods that rely on fixed view grids or dense multi-camera capture, PVSNet operates on arbitrary camera positions and orientations, making it suitable for unconstrained, real-time scenarios.
+The ability to embed full 6DoF poses directly into the synthesis pipeline allows our method to generalize across both panoramic and planar input modalities without requiring view-specific encodings or ray models.
+2.2. Single-View Input
+Single-view input methods aim to synthesize novel views from just one image, primarily addressing the scarcity of limited input data.
+Many of these techniques employ deep learning models to infer depth or scene structure from a single image.
+For instance, PixelNeRF (Yu et al., 2021) and VisionNeRF (Lin et al., 2023) generalize NeRF to new scenes by conditioning on image features at large scale.
+Although the PixelNeRF and VisionNeRF show promising results, they face challenges such as scale ambiguity and large baseline camera movements, which degrade rendering quality in more extreme viewpoints.
+To mitigate issues like scale ambiguity and substantial viewpoint changes, diffusion-based approaches have emerged.
+For example, 3DiM (Watson et al., 2022) employs pose guided diffusion to improve performance, particularly in scenarios with large translational movement.
+Fine-tuning diffusion models on large-scale synthetic datasets (Deitke et al., 2023) has also shown potential (Liu et al., 2023, 2024) in handling the scale ambiguity, although these methods are hindered by slow inference times, and highest memory footprint compared to all other methods discussed so far.
+Other single-view input methods such as 3D Photo (Shih et al., 2020) and SLIDE (Jampani et al., 2021) focus on inferring views through learned layered depth inpainting.
+These methods can yield visually compelling results, especially for large baselines, but often require RGB+D input (Shih et al., 2020) or rely on multi-stage pipelines with monocular depth estimation, layering, and hole-filling (Jampani et al., 2021), making them unsuitable for real-time applications.
+As an alternative to diffusion-based frameworks, multiplane image (MPI) representations approximate a scene using a series of parallel planes at fixed depths, each containing a color and alpha map.
+This representation allows view synthesis based on a fixed (Zhou et al., 2018; Flynn et al., 2019; Tucker and Snavely, 2020; Li and Kalantari, 2020) or variable (Li and Kalantari, 2020) number of planes.
 While MPI-based methods achieve superior visual quality, their computational demands remain high, especially during MPI layer prediction.
-Large-scale GANs (Koh et al., 2023) have also been explored for their ability to render fast and handle large baselines, though they introduce significant memory overhead, limiting their practicality.
 More recent innovations, such as NViST (Jang and Agapito, 2024), incorporate vision transformers (ViTs) (Dosovitskiy, 2020) into view synthesis architectures, demonstrating strong generalization capabilities in real-world settings.
-By leveraging masked autoencoders (MAEs) and adaptive layer normalization conditioned on camera parameters, NViST delivers high-quality results, even for out-of-distribution scenes.
-Our proposed method, PLFNet+ (described in Sec 4), contributes to the learning-based category by extending the views from a single input image while diverging from diffusion and transformer-based methods.
-By utilizing a lightweight design, PLFNet+ reduces computational overhead, enabling real-time rendering.
-This efficiency makes it particularly suited for dynamic environments where fast processing is crucial, addressing key challenges in the field of view synthesis.
-3. Method: Light Field Reconstruction - PLFNet
-We first define a Position Aware Light Field Reconstruction network called ”PLFNet” which aims to reconstruct the LF from a single input image by iterative reconstructing individual sub-aperture images.
-The source image serves as the basis for reconstructing the surrounding LF , where x represents the angular coordinate , and signifies the center of the LF.
-Each sub-aperture image at location , specified as , undergoes reconstruction via:
-This two stage pipeline has been shown in Figure 1
+By leveraging masked autoencoders (MAEs) and adaptive layer normalization conditioned on camera parameters, NViST delivers high-quality results, even for out-of-distribution views.
+Despite these strengths, the reliance of NViST on a ViT backend and patch-wise image processing leads to slower inference at higher resolutions.
+Meanwhile, large-scale GANs (Koh et al., 2023) offer rapid rendering and handle large baselines effectively at high resolution from a single ODI, but it also incurs significant memory overhead, limiting the practical usage in resource-constrained settings.
+Another relevant line of work is MegaParallax (Bertel et al., 2019), which aims to produce motion parallax in ODI using layered mesh generated from monocular depth maps and image warping.
+While it succeeds in synthesizing novel viewpoints from a single ODI, the pipeline involves explicit depth estimation, foreground segmentation, and layered rendering, which introduce computational overhead and make real-time operation impossible.
+In summary, despite the advances, single-view methods frequently grapple with achieving both real-time performance and high-quality novel view generation in live applications.
+Addressing this balance between efficiency and fidelity remains a critical goal for immersive telepresence and related applications.
+The limitations identified here, which includes scale ambiguity, heavy computational demands, and memory constraints, inform the design of our proposed approach PVSNet (described in Section 3), which aims to reconcile the trade-offs these methods face.
+3. Method: Position-Aware View-Synthesis
+PVSNet is a two stage model which, given the source image and the target position , aims to synthesize the novel image at the corresponding target viewpoint.
+Our approach consists of two key stages as shown in Figure 2.
+First, the target position is embedded into a high-dimensional space through positional embedding which is discussed in Section 3.1, and second, new views are rendered from these positions, which is detailed in Section 3.2.
 3.1. Positional Embedding
-We define the target position by the coordinate vector , where , and represent translations associated with angular coordinates.
-Given limited baseline shifts of LF reconstruction, directly using normalized values of can suffice for training the network on discrete locations but querying for any continuous values of .
-Coordinate Normalization: The target coordinates are normalized into to ensure consistency across varying ranges of , . The normalization process is given by
+We aim to synthesize views from a single input image and a target pose , defined by 3D translation and rotation representing yaw, pitch, and roll, respectively.
+While recent approaches such as Cameras as Rays (Zhang et al., 2024) propose learning patch wise ray bundles for pose estimation, these are designed for multi-view image setups and rely on ray supervision and diffusion sampling.
+In contrast, our goal is real-time view synthesis from a single image, where camera pose is already known.
+Thus, we adopt a compact 6D vector representation and learn a high-dimensional embedding to support feed-forward synthesis at high frame rates.
+Although direct regression using may suffice for small viewpoint changes, it becomes insufficient for larger baselines.
+To better capture pose dependent variations, we employ a positional embedding strategy that combines positional encoding, inspired by NeRF (Mildenhall et al., 2021), along with our proposed MLP module for enhanced representation of the target coordinate .
+Coordinate Normalization: First, the target coordinates denoted by are normalized into to ensure consistency across varying ranges of viewpoints in scenes and training samples :
+where denote the minimum and maximum values across the training set for each component of .
+This ensures all elements of the normalized pose vector .
 This normalization step ensures that the coordinates lie within a fixed range, which is crucial for subsequent encoding.
-Input Channel: To incorporate the normalized coordinates into the network, the values of are expanded into a tensor , where the first channel holds the values, and the second channel holds the values
-This tensor has the same spatial resolution as the input image, ensuring that positional information is aligned with the image features spatially.
+Positional Encoding: To project the normalized coordinates into a higher-dimensional space, we apply positional encoding.
+This involves concatenating multiple sine and cosine functions of varying frequencies
+where is the normalized 6D coordinate vector, represents the encoding depth, and .
+The parameter is a scaling constant based on the domain of .
+For our experiments, we use , resulting in an output size of dimensions, which provides enhanced positional awareness (see Table 5 for ablation study results).
+Learning Based Reprojection: Once the positional encoding has been computed, we further enhance the representation by passing it through an MLP.
+This MLP maps the given encoded vector into a much higher-dimensional space, which has proven to work well with scene regression task as shown in NeRF (Mildenhall et al., 2021).
+This process allows for a more expressive and flexible feature space which, unlike NeRF, can be adapted to different scenes since we only consider the coordinates in this representation and not the scene properties at this point.
+Our approach is similar to the adaptive layer normalization techniques used in NViST (Jang and Agapito, 2024).
+Let be the output of the positional encoding, where is the dimensionality of the encoding. The MLP projects this vector into a high-dimensional space , where is chosen based on the desired resolution:
+Then, the vector is reshaped into a feature map of size , where , w:h has the same aspect ratio as the original image, and .
+This reshaping operation can be expressed using the inverse vectorization notation as
+Finally, the feature map is upsampled using nearest neighbor method to match the spatial resolution of the input image, denoted by , thus we set .
+This upsampling is performed using a scaling factor , as follows
+The final upsampled representation retains the positional information of the target coordinate and is subsequently used by the rendering network to generate the novel view.
 3.2. Rendering Network
-The role of the rendering network in PLFNet is to synthesize a novel target image from the input image and the positional embedding , obtained from the positional embedding stage.
+The role of the rendering network in PVSNet is to synthesize a target image from the input image and the positional embedding , obtained from the positional embedding stage.
 Our network architecture is inspired by LFSphereNet (Gond et al., 2023) and consists of two distinct encoder branches followed by a decoder network.
@@
 Unlike the first encoder, which only processes the image, focuses on refining the feature maps based on the target position.
+This is done by capturing the high‐frequency variations in geometry and appearance much more effectively.
 We first concatenate the input image and the positional embedding along the channel dimension, resulting in a stacked input tensor
-where is the number of channels in . In our case, for LF Reconstruction.
+where is the number of channels in . In our case, .
 This stacked tensor is passed through a series of convolutional layers that preserve the spatial dimensions, ensuring pixel-level positional information is retained throughout the encoding process, as demonstrated by LFSphereNet (Gond et al., 2023).
-The encoder produces a feature map , where is the number of feature channels, and skip connections are generated from intermediate layers for later use in the decoder.
+In the last two layers, Max Pooling is employed to downsample the feature maps to match the spatial dimension of feature map .
+The encoder produces a feature map , where is the number of feature channels.
+Additionally, skip connections are generated from intermediate layers for later use in the decoder.
 Formally, this can be expressed as
@@
 3.3. Training Loss
-To optimize our PLFNet, we employ a combination of the L1 loss and the Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003) loss.
-This combination leverages the complementary strengths of each loss function.
-The L1 loss minimizes pixel-wise differences between the predicted sub-aperture images and the ground truth images , which helps to maintain accurate reconstruction of intensity values.
-The MS-SSIM loss, on the other hand, is a perceptually motivated metric that focuses on preserving structural information and texture across multiple scales, which correlates better with human visual perception as shown in image restoration task (Zhao et al., 2016).
-The total loss function is defined as a weighted sum of the L1 and MS-SSIM losses
-where is a weighting factor that balances the contribution of the two losses.
-This combined loss function enables the network to generate high-quality light field reconstructions that are both pixel-accurate and perceptually consistent.
-3.4. Implementation: PLFNet
-The input of PLFNet is a single RGB image and the target position .
-Here and which are similar to the spatial resolution of input image .
-We set because the datasets (Srinivasan et al., 2017; Raj et al., 2016; Rerabek and Ebrahimi, 2016) are organized as a 2D grid, with only two angular coordinates considered.
-PLFNet was implemented in PyTorch and trained with the following hyperparameters: batch size of 24, learning rate of 0.003, adam optimizer, MS-SSIM and L1 loss.
+To optimize our PVSNet, we combine four loss functions: L1 loss, Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003) loss, Focal Frequency Loss (FFL) (Jiang et al., 2021) and perceptual loss (Chen and Koltun, 2017) .
+Each of these losses contributes to different aspects of the synthesis process, improving both spatial and frequency domain accuracy.
+L1 Loss and MS-SSIM Loss are combined to capture fine-grained details, especially when dealing with larger baselines.
+The L1 loss minimizes pixel-wise differences between the predicted novel view image and the ground truth image , ensuring accurate reconstruction of intensity values.
+On the other hand, MS-SSIM is a perceptually motivated metric that focuses on preserving structural information and texture at multiple scales, correlating better with human visual perception, as demonstrated in image restoration tasks (Zhao et al., 2016).
+Focal Frequency Loss (FFL) is utilized to enhance the quality of the synthesized views, particularly in the frequency domain.
+We incorporate Focal Frequency Loss (FFL) (Jiang et al., 2021) to address spectral bias in neural networks by focusing on the frequency components that are hardest to synthesize.
+This helps the model to improve the reconstruction of high-frequency details, which are crucial for generating realistic novel views.
+The FFL is defined as:
+where and represent the real and predicted frequency components at position in the Fourier domain, and is a dynamic weight that emphasizes the hardest-to-synthesize frequencies.
+Perceptual Loss is used to improve the perceptual quality of the synthesized views by comparing high-level feature activations between the predicted and ground truth images.
+Inspired by (Chen and Koltun, 2017), we extract intermediate feature maps from a pretrained VGG-19 network and minimize the L1 difference between corresponding layers.
+These layers represent increasing levels of abstraction from edges to object parts and thus guide the synthesis model to preserve both fine grained detail and global structure.
+Specifically, we define the perceptual loss as:
+where denotes the activation from layer of the VGG-19 network, and is a weighting factor inversely proportional to the number of elements in that layer.
+This perceptual supervision helps produce sharper, more semantically coherent outputs, particularly in areas where pixel-wise losses are insufficient to capture visual fidelity.
+The total loss function for PVSNet is then defined as:
+where is the weighting factor that balances the contribution of L1 and MS-SSIM losses and is a weight factor that balances the contribution of the FFL relative to the L1 and MS-SSIM losses.
+This combined loss function enables PVSNet to generate sharper and more realistic novel views by optimizing for both spatial accuracy and high-frequency details.
+3.4. Implementation
+The input of PVSNet is a single RGB image and the target position , for the positional embedding output we set since we use a single channel positional map.
+PVSNet was implemented in PyTorch111Code and checkpoint for our model will be made available upon acceptance of this paper at https://github.com/Realistic3D-MIUN/PVSNet , and trained with the following hyperparameters: batch size of 24, learning rate of 0.003, adam optimizer, MS-SSIM, L1 and FFL loss.
 After each 30 epochs, the learning rate was decreased by a factor of 0.2.
@@
 The training duration was 48 hrs.
-For the dataset split, 75% of the data was used for training, 12.5% for validation and 12.5% for testing.
-4. Method: View Synthesis - PLFNet+
-Our approach for view synthesis builds upon PLFNet by extending the positional embedding to handle wider baselines with translation along the -axis.
-Here, represents the synthesized image corresponding to the target viewpoint.
-Our approach consists of two key stages as shown in Figure 2: first, the target position is embedded into a high-dimensional space through positional embedding (discussed in Section 4.1), and second, novel views are rendered from these positions, which is detailed in Section 4.2.
-4.1. Extended Positional Embedding
-We define the target position by the coordinate vector , where , , and represent translations from the input image .
-As the baseline increases and the search space extends along the axis, it becomes necessary to project the target coordinates into a higher-dimensional space to accurately capture positional information.
-To address this, we employ a positional embedding strategy that combines positional encoding, inspired by NeRF (Mildenhall et al., 2021), along with a MLP for enhanced representation of the target coordinate .
-Coordinate Normalization: Similar to the normalization strategy in Section 3.1, we convert the target coordinates into their normalized form .
-Positional Encoding: To project the normalized coordinates into a higher-dimensional space, we apply positional encoding.
-This involves concatenating multiple sine and cosine functions of varying frequencies
-where is the normalized 3D coordinate vector, represents the encoding depth, and .
-The parameter is a scaling constant based on the domain of .
-For our experiments, we use , therefore yields an output of size dimensions, which provides enhanced positional awareness (see Table 6 for ablation study results).
-Learning Based Reprojection: Once the positional encoding has been computed, we further enhance the representation by passing it through an MLP.
-This MLP maps the low-dimensional encoded vector into a higher-dimensional space, and it has proven to work well with scene regression task as shown in NeRF (Mildenhall et al., 2021).
-This process allows for a more expressive and flexible feature space, and our approach is similar to the adaptive layer normalization techniques used in NViST (Jang and Agapito, 2024).
-Let be the output of the positional encoding, where is the dimensionality of the encoding.
-The MLP projects this vector into a high-dimensional space , where is chosen based on the desired resolution.
-Mathematically, the projection is represented as
-After this, the vector is reshaped into a feature map of size , where . This reshaping operation can be expressed using the inverse vectorization notation as
-Finally, the feature map is upsampled to match the spatial resolution of the input image, denoted by . This upsampling is performed using a scaling factor , as follows
-The final upsampled representation retains the positional information of the target coordinate and is subsequently used by the rendering network to generate the novel view.
-4.2. Rendering Network
-The rendering network remains similar to Section 3.2. However, in the case of view synthesis, the positional embedding is denoted by a single channel, simplifying the input to the second encoder as
-The output of the rendering network is the target image .
-4.3. Training Loss
-For the view synthesis task, we extend the loss function previously defined in Section 3.3 to better capture fine-grained details, especially with larger baselines.
-We maintain the same combination of L1 loss and MS-SSIM loss, which focuses on minimizing pixel-wise differences and preserving structural consistency, respectively.
-However, to further enhance the quality of the synthesized views, particularly in the frequency domain, we incorporate the Focal Frequency Loss (FFL) (Jiang et al., 2021) into the total loss function.
-The FFL addresses the spectral bias in neural networks by adaptively focusing on the frequency components that are hardest to synthesize.
-This allows the model to improve the reconstruction of high-frequency details, which are critical for realistic novel view generation.
-The FFL is defined as
-where and represent the real and predicted frequency components at position in the Fourier domain, and is a dynamic spectrum weight that emphasizes hard-to-synthesize frequencies.
-Thus, the total loss function for PLFNet+ is updated to
-where is the Focal Frequency Loss, and is a weight balancing its contribution relative to the L1 and MS-SSIM losses.
-This augmented loss function allows PLFNet+ to generate sharper, more realistic novel views by not only optimizing for spatial accuracy but also addressing frequency domain discrepancies.
-4.4. Implementation: PLFNet+
-We keep the same implementation as in PLFNet (described in Section 3.4, but set , .
-We also set as the positional maps is single channel.
-Due to limitation of training data when using our synthetic dataset we employ multi stage training strategies to reach satisfactory results in training process.
+Due to limitation of training data when using our synthetic dataset we employ multi stage training strategies to reach satisfactory results in the training process.
 3.4.1. Multi Stage Training
@@
 Stage 1 serves as a preconditioning step, training the network on a large-scale light field dataset (Srinivasan et al., 2017).
-While the baseline between the views in each light field is relatively narrow, the dataset’s large scale enables the network to learn both high- and low-level image features, implicitly understanding depth.
-In this stage, the network is conditioned to synthesize novel views within a constrained light field volume, similar to the reconstruction process described in PLFNet (Section 3).
+While the baseline between the views in each light field is relatively narrow, the large scale of dataset enables the network to learn both high and low level image features, implicitly understanding depth.
+In this stage, the network is conditioned to synthesize novel views within a constrained light field volume.
 Once the network demonstrates proficiency in this task, we transition to the next stage.
@@
 Following this, the network is prepared for the final stage of training, which focuses on the joint optimization of the MLP and rendering network with more random positions.
-Stage 3 involves training the network on a synthetic dataset containing random camera viewpoints within an cubical volume.
+Stage 3 involves training the network on our synthetic dataset (discussed in Section 4.1.1) containing random camera viewpoints within a cubical volume.
 The increase in the spatial search space necessitates the use of the positional embedding module for accurate view synthesis.
 In this stage, the network is fine-tuned to handle more complex camera translations and larger scene variations, fully utilizing the positional embedding to enhance rendering precision.
+| Classroom | Italian Flat |
+| Lone Monk | Barbershop |
 4. Experimental Evaluation
-In this section, we present a comprehensive evaluation of our proposed PLFNet and PLFNet+ architectures.
-We begin by assessing the performance of PLFNet+ for view synthesis on datasets with varying baselines and complexities.
-Following this, we demonstrate the capability of PLFNet in reconstructing full light fields from a single input image.
+In this section, we present a comprehensive evaluation of our proposed PVSNet architecture.
+We begin by assessing the performance of PVSNet for view synthesis on datasets with varying baselines and complexities.
+Following this, we demonstrate the capability of PVSNet for small baseline light field reconstruction and view synthesis within LF baseline from a single input image.
 The experiments for view synthesis are detailed in Section 4.1, while light field reconstruction results are discussed in Section 4.2.
-5.1. View Synthesis
 | Dataset | Method | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
@@
 | 256x256 | AdaMPI (Han et al., 2022) | 24.9252 | 0.8092 | 0.9145 | 0.6720 | 0.0680 | 0.0455 |
-| PLFNet+ (Ours) | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 | |
+| PVSNet (Ours) | 31.2258 | 0.9386 | 0.9793 | 0.8702 | 0.0613 | 0.0275 | |
+| PVSNet-Lite (Ours) | 28.7894 | 0.8902 | 0.9577 | 0.7655 | 0.0942 | 0.0566 | |
 | SinMPI (Pu et al., 2023) | 18.3441 | 0.4599 | 0.6885 | 0.3621 | 0.2251 | 0.3966 | |
@@
 | 512x512 | AdaMPI (Han et al., 2022) | 23.1416 | 0.7510 | 0.8566 | 0.4353 | 0.0751 | 0.0770 |
-| PLFNet+ (Ours) | 26.3339 | 0.8334 | 0.9307 | 0.5633 | 0.2191 | 0.1947 | |
-| SinMPI (Pu et al., 2023) | 14.5548 | 0.4271 | 0.6621 | 0.3706 | 0.2867 | 0.4024 | |
-| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 17.1759 | 0.5454 | 0.7314 | 0.3819 | 0.2028 | 0.2415 |
-| 256x256 | AdaMPI (Han et al., 2022) | 23.6155 | 0.8263 | 0.9345 | 0.7232 | 0.1055 | 0.0797 |
-| PLFNet+ (Ours) | 22.8933 | 0.7608 | 0.9070 | 0.5534 | 0.2378 | 0.2560 | |
-| SinMPI (Pu et al., 2023) | 15.7272 | 0.4740 | 0.6712 | 0.3095 | 0.2392 | 0.3775 | |
-| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 18.1375 | 0.5444 | 0.7499 | 0.2906 | 0.1750 | 0.2758 |
-| 512x512 | AdaMPI (Han et al., 2022) | 21.5931 | 0.7425 | 0.8925 | 0.5215 | 0.1156 | 0.1559 |
-| PLFNet+ (Ours) | 20.2596 | 0.6229 | 0.8001 | 0.3007 | 0.2758 | 0.4226 |
+| PVSNet (Ours) | 30.0085 | 0.9254 | 0.9733 | 0.8408 | 0.0733 | 0.0329 | |
+| PVSNet-Lite (Ours) | 25.8741 | 0.8135 | 0.9100 | 0.6063 | 0.0852 | 0.0778 | |
+| SinMPI (Pu et al., 2023) | 14.8544 | 0.4669 | 0.6484 | 0.3521 | 0.2940 | 0.5111 | |
+| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 17.6991 | 0.5822 | 0.7411 | 0.4032 | 0.2161 | 0.2490 |
+| 256x256 | AdaMPI (Han et al., 2022) | 22.1027 | 0.8093 | 0.9241 | 0.7180 | 0.1152 | 0.0849 |
+| PVSNet (Ours) | 24.0435 | 0.8462 | 0.9445 | 0.7518 | 0.0825 | 0.0600 | |
+| PVSNet-Lite (Ours) | 21.3391 | 0.7555 | 0.9122 | 0.6195 | 0.1147 | 0.0915 | |
+| SinMPI (Pu et al., 2023) | 15.1025 | 0.4561 | 0.6500 | 0.3149 | 0.2250 | 0.3555 | |
+| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 18.2453 | 0.6290 | 0.7851 | 0.3003 | 0.1750 | 0.2711 |
+| 512x512 | AdaMPI (Han et al., 2022) | 21.3099 | 0.7356 | 0.8792 | 0.5099 | 0.1196 | 0.1673 |
+| PVSNet (Ours) | 22.8850 | 0.7764 | 0.9015 | 0.6095 | 0.1070 | 0.1437 | |
+| PVSNet-Lite (Ours) | 19.8910 | 0.6200 | 0.7842 | 0.4613 | 0.1364 | 0.1894 |
 | Method | 256x256 | 512x512 |
@@
 | AdaMPI (Han et al., 2022) | 0.2261 / 4 | 0.3254 / 3 |
-| PLFNet+ (Ours) | 0.0074 / 135 | 0.0086 / 116 |
+| PVSNet (Ours) | 0.0161 / 62 | 0.0181 / 55 |
+| PVSNet-Lite (Ours) | 0.0081 / 123 | 0.0096 / 104 |
+4.1. View Synthesis
 4.1.1. Datasets
-To evaluate the performance of PLFNet+ on view synthesis, we used two datasets.
-The first is a custom Blender dataset, created specifically for this task, which contains images from different scenes with camera movements constrained within a 20 cm cubical volume.
-The second dataset is COCO (Caesar et al., 2018), which we use during training in the same manner as AdaMPI (Han et al., 2022), where depth-image-based rendering (DIBR) with hole filling is employed to generate the training data.
+To evaluate the performance of PVSNet on view synthesis, we use two datasets: a custom Blender dataset tailored for this task and the COCO dataset (Caesar et al., 2018) for more diverse scenes to demonstrate broader generalizability.
+Blender dataset was created by selecting four freely available virtual scenes from the Blender demo website—namely Lone Monk, Classroom, Barbershop, and Italian Flat as shown in Figure 4.
+For each scene, we first positioned the camera at various initial locations and then applied random camera movements constrained within a cubical volume around each initial position.
+These random movements allowed us to capture multiple images per scene, all within the specified region.
+The resulting set of images from these scenes constitutes our custom Blender dataset222Dataset url will be made available upon acceptance of this paper.
+COCO dataset was adopted with a similar approach to AdaMPI (Han et al., 2022) for generating training data, which involved utilizing depth-image-based rendering (DIBR) with hole filling.
 During evaluation, we take an input image and synthesize a novel image at a specified target location.
 The synthesized image is then treated as input to generate an image back at the original location.
-This process enables us to compare the reconstructed image with the ground truth, which is the original input image, allowing for an assessment of rendering quality.
-Both datasets are evaluated at resolutions of and and a movement within a volume of .
+This process enables us to compare the reconstructed image with the ground truth (the original input image), thereby allowing for an assessment of rendering quality.
+Both datasets are evaluated at resolutions of and .
 4.1.2. Evaluation Metrics
-To objectively evaluate the quality of the synthesized novel views, we employed multiple standard and advanced metrics.
+To objectively evaluate the quality of the synthesized views, we employed multiple standard and advanced metrics.
 These include peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) (Wang et al., 2004), as well as multi-scale structural similarity index measure (MS-SSIM) (Wang et al., 2003), pixel- based visual information fidelity (VIFP) (Sheikh and Bovik, 2006), deep image structure and texture similarity (DISTS) (Ding et al., 2020), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018).
@@
 4.1.3. Results
-Quantitative Results: The quantitative results for view synthesis across different datasets and resolutions are presented in Table 1.
-Our proposed PLFNet+ consistently outperforms AdaMPI (Han et al., 2022) on the Blender dataset and achieves competitive results on the COCO (Caesar et al., 2018) dataset.
-Although AdaMPI demonstrates slightly better performance in perceptual metrics such as DISTS and LPIPS, our method remains competitive across a range of metrics, especially on PSNR and SSIM.
-Interestingly, we observe a slight performance drop in AdaMPI and PLFNet+ when increasing the resolution from to .
+In this section, we compare the view synthesis performance of PVSNet with state-of-the-art methods, namely AdaMPI (Han et al., 2022), TMPI (Khan et al., 2023), and SinMPI (Pu et al., 2023).
+To ensure a fair comparison, all methods were trained on the same dataset, using their respective original implementations and training strategies.
+Additionally, we introduce a lightweight variant called PVSNet-Lite, specifically optimized for maximum inference speed with minimal quality degradation.
+This version uses a simplified rendering network and reduced feature dimensionality to significantly boost FPS, making it ideal for latency-critical applications.
+The details of this PVSNet-Lite will be available in our open sourced code11footnotemark: 1.
+Quantitative Image Quality: The quantitative results for view synthesis across different datasets and resolutions are presented in Table 1.
+Our proposed PVSNet consistently outperforms AdaMPI (Han et al., 2022) on the Blender dataset and achieves competitive results on the COCO (Caesar et al., 2018) dataset.
+Our method remains competitive across a range of metrics, especially on PSNR and SSIM.
+Interestingly, we observe a slight performance drop in AdaMPI and PVSNet when increasing the resolution from to .
 This highlights the challenges of scaling models without increasing their capacity to handle higher-resolution inputs.
-In contrast to other methods that explicitly rely on depth estimation using the DPT model, our approach learns scene depth implicitly from RGB input alone, demonstrating the flexibility of our method without the need for pre-computed depth maps.
-When evaluating inference time, PLFNet+ excels, particularly at higher resolutions.
-As shown in Table 2, our method achieves an inference time of 0.0086 seconds for a resolution of , outperforming all other methods by a significant margin.
-This makes PLFNet+ highly suitable for real-time applications, where rapid generation of novel views is critical, such as in dynamic video streams or interactive virtual environments.
-Qualitative Results: Figure 4 provides a visual comparison of the view synthesis results for different methods on the COCO dataset.
+In contrast to other methods that explicitly rely on depth estimation using the DPT (Ranftl et al., 2022) model, our approach learns scene depth implicitly from RGB input alone, demonstrating the flexibility of our method without the need for pre-computed depth maps.
+Quantitative Processing Time: To better illustrate the trade off between inference speed and image quality, we present a comparison of all methods in Figure 5.
+The plot shows MS-SSIM versus FPS at both and resolutions, clearly highlighting that PVSNet achieves an optimal balance, while PVSNet-Lite pushes the boundary further by significantly improving FPS with only a modest drop in quality.
+As shown in Table 2, our method achieves an inference time of 0.0181 seconds (55 fps) for a resolution of , outperforming all other methods by a factor 18 .
+We also report results for PVSNet-Lite in Table 2, which achieves even higher frame rates (around 104 FPS at ) with only a marginal drop in visual quality.
+This makes PVSNet and PVSNet-Lite highly suitable for real-time applications, where rapid generation of novel views is critical, such as in dynamic video streams or interactive virtual environments.
+A pre-recorded video in our supplementary materials shows the performance in dynamic video streams.
+Qualitative Image Quality: Figure 6 provides a visual comparison of the view synthesis results for different methods on the COCO dataset.
 Each method’s output is accompanied by error maps, allowing for a more detailed analysis of their accuracy.
-SinMPI (Pu et al., 2023) struggles with rendering artifacts, particularly in regions with complex textures and geometry, as evidenced by the error maps.
-These issues are likely a result of its depth estimation and scene representation.
-TMPI (Khan et al., 2023) also displays notable tiling artifacts due to its approach of partitioning the scene into multiple planes, leading to visible seams and discontinuities.
-In contrast, both AdaMPI and our proposed PLFNet+ produce the most visually accurate results, with minimal artifacts.
-AdaMPI shows slightly better performance in some scenarios, particularly in handling fine details, but our method’s ability to efficiently handle reflections, occlusions, and fine textures stands out.
-While AdaMPI may have a slight edge in perceptual quality, our method’s superior inference speed makes it the more practical choice for real-time applications, balancing high-quality synthesis with the demands of time-sensitive environments.
-5.2. Extending View Synthesis: Light Field Reconstruction
-5.2.1. Datasets
-For LF reconstruction, we evaluated our method using three publicly available datasets: Lytro Flowers (Srinivasan et al., 2017), Stanford Light Field Archive (Raj et al., 2016), and JPEG-Pleno (Rerabek and Ebrahimi, 2016).
-These datasets are widely used for benchmarking light field reconstruction algorithms.
-The extracted light field images had dimensions of 352x512 with a 7x7 angular resolution.
-We trained PLFNet on 75% of the Lytro Flowers and Stanford datasets, with the remaining 25% reserved for testing.
-The JPEG-Pleno dataset, however, was exclusively used for testing to evaluate the generalization capability of our model, as it was not included in the training process.
-5.2.2. Evaluation Metrics
-We utilize similar evaluation metrics as described in Section 5.1.2.
-Since the light field reconstruction works in very narrow baseline between each sub-aperture image, and has discrete camera locations, the search space is much smaller therefore we get much better scores in these metrics.
 | Dataset | Method | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
@@
 | LFSphereNet | 41.3719 | 0.9461 | 0.9868 | 0.9060 | 0.0812 | 0.0512 | |
-| PLFNet | 37.2606 | 0.9601 | 0.9914 | 0.9062 | 0.0608 | 0.0229 | |
+| PVSNet | 37.2606 | 0.9601 | 0.9914 | 0.9062 | 0.0608 | 0.0229 | |
 | Stanford (Raj et al., 2016) | NoisyLFRecon | 35.4917 | 0.9572 | 0.9771 | 0.8380 | 0.0718 | 0.0395 |
@@
 | LFSphereNet | 40.9830 | 0.9488 | 0.9797 | 0.8842 | 0.0556 | 0.0300 | |
-| PLFNet | 36.8530 | 0.9601 | 0.9927 | 0.8679 | 0.0553 | 0.0182 | |
+| PVSNet | 36.8530 | 0.9601 | 0.9927 | 0.8679 | 0.0553 | 0.0182 | |
 | JPEG Pleno (Rerabek and Ebrahimi, 2016) | NoisyLFRecon | 36.6275 | 0.9575 | 0.9687 | 0.8549 | 0.0718 | 0.0435 |
@@
 | LFSphereNet | 39.2624 | 0.9429 | 0.9879 | 0.9069 | 0.0737 | 0.0425 | |
-| PLFNet | 33.8982 | 0.9106 | 0.9838 | 0.8683 | 0.0590 | 0.0261 |
-| Method | Purpose | Inference Time (s) |
-| NoisyLFRecon | LF reconstruction | 1.1830 |
-| DGLF | LF reconstruction | 1.5009 |
-| DALF | LF reconstruction | 0.0583 |
-| IR-V | LF reconstruction | 0.1429 |
-| LFSphereNet | LF reconstruction | 0.0008 |
-| PLFNet | Novel view synthesis | 0.0980 |
-| NoisyLFRecon | DGLF | DALF | IR-V | LFSphereNet | PLFNet |
+| PVSNet | 33.8982 | 0.9106 | 0.9838 | 0.8683 | 0.0590 | 0.0261 |
+SinMPI (Pu et al., 2023) struggles with rendering artifacts, particularly in regions with complex textures and geometry, as evidenced by the error maps.
+These issues are likely a result of its depth estimation and scene representation.
+TMPI (Khan et al., 2023) also displays notable tiling artifacts due to its approach of partitioning the scene into multiple planes, leading to visible seams and discontinuities.
+In contrast, both AdaMPI and our proposed PVSNet produce the most visually accurate results, with minimal artifacts.
+Our method’s ability to efficiently handle reflections, occlusions, and fine textures stands out, in addition the inference speed makes it the more practical choice for real-time applications, balancing high-quality synthesis with the demands of time-sensitive environments.
+Additional visual results when running the PVSNet on in-the-wild photos are presented in our supplementary materials video.
+4.2. Light Field Reconstruction and View Synthesis
+In this section, we demonstrate alternative approach for view synthesis where Light Field reconstruction is the intermediate step.
+We focus on capability of proposed PVSNet architecture in reconstructing full light fields from a single input image.
+In this case study, we target very small baseline of the LFs, and we use a smaller version of PVSNet by modifying the positional embedding step to only use the coordinate normalization during training and inference.
+The setup of PVSNet which utilizes positional embedding with only the coordinate normalization has been studied in detail in our ablation study in Section 5.1, where we show that this setup is sufficient to handle small baselines.
+4.2.1. Datasets
+For LF reconstruction, we evaluated our method using three publicly available datasets: Lytro Flowers (Srinivasan et al., 2017), Stanford Light Field Archive (Raj et al., 2016), and JPEG-Pleno (Rerabek and Ebrahimi, 2016).
+These datasets are widely used for benchmarking light field reconstruction algorithms.
+The extracted light field images had dimensions of 352x512 with a 7x7 angular resolution.
+We trained PVSNet on 75% of the Lytro Flowers and Stanford datasets, with the remaining 25% reserved for testing.
+The JPEG-Pleno dataset, however, was exclusively used for testing to evaluate the generalization capability of our model, as it was not included in the training process.
+4.2.2. Evaluation Metrics
+We utilize similar evaluation metrics as described in Section 4.1.2.
+Since the light field reconstruction works in very narrow baseline between each sub-aperture image, and has discrete camera locations, the search space is much smaller.
+Therefore we expect much better scores in these metrics.
+| Method | Purpose | LF Reconstruction (sec) | View Synthesis (sec) |
+| NoisyLFRecon | LF reconstruction | 1.0923 | 2.3933 |
+| DGLF | LF reconstruction | 1.3062 | 2.6072 |
+| DALF | LF reconstruction | 0.0577 | 1.3587 |
+| IR-V | LF reconstruction | 0.1342 | 1.4352 |
+| LFSphereNet | LF reconstruction | 0.0008 | 1.3018 |
+| PVSNet | View synthesis | 0.0850 | 0.0079 |
+| NoisyLFRecon | DGLF | DALF | IR-V | LFSphereNet | PVSNet |
 | (Zhou et al., 2021) | (Zhou et al., 2020) | (Cun et al., 2019) | (Han and Xiang, 2022) | (Gond et al., 2023) | (Ours) |
 4.2.3. Results
-Quantitative Results: The quantitative results of our light field reconstruction experiments are presented in Table 3.
-Our proposed PLFNet demonstrates competitive performance compared to state-of-the-art methods, including NoisyLFRecon (Zhou et al., 2021), DGLF (Zhou et al., 2020), DALF (Cun et al., 2019), IR-V (Han and Xiang, 2022), and LFSphereNet (Gond et al., 2023).
+Quantitative Image Quality: The quantitative results of our light field reconstruction experiments are presented in Table 3.
+Our proposed PVSNet demonstrates competitive performance compared to state-of-the-art methods, including NoisyLFRecon (Zhou et al., 2021), DGLF (Zhou et al., 2020), DALF (Cun et al., 2019), IR-V (Han and Xiang, 2022), and LFSphereNet (Gond et al., 2023).
 All models were trained from scratch using identical datasets for a fair comparison.
@@
 (2021, 2020) and Han and Xiang (2022), while DALF (Cun et al., 2019) was re-implemented in PyTorch according to the original paper.
-Although DGLF and LFSphereNet generate the full light field in a single pass, both DALF and our PLFNet reconstruct each sub-aperture image iteratively.
-For the Stanford dataset, which offers fewer training samples compared to Lytro Flowers, LFSphereNet delivered the best results across most metrics, with PLFNet showing competitive performance, particularly excelling in SSIM and MS-SSIM.
-In terms of computational efficiency, PLFNet demonstrated significant improvements as shown in Table 4.
-Our PLFNet achieved an inference time of 0.0980 seconds for whole LF reconstruction which included forward passes.
-However, per view inference on average is around 0.006 seconds highlighting its potential for real-time applications.
-Qualitative Results: The qualitative evaluation of light field reconstruction is illustrated in Figure 5.
-We compare the visual performance of our proposed PLFNet against state-of-the-art methods for light field views, focusing on the top right corner sub-aperture view.
+Although DGLF and LFSphereNet generate the full light field in a single pass, both DALF and our PVSNet reconstruct each sub-aperture image iteratively.
+For the Stanford dataset, which offers fewer training samples compared to Lytro Flowers, LFSphereNet delivered the best results across most metrics, with PVSNet showing competitive performance, particularly excelling in SSIM and MS-SSIM.
+However, we note that while LFSphereNet and the other methods in this comparison were specifically designed for light field reconstruction, PVSNet is optimized for view synthesis within the narrow baseline of light fields.
+Consequently, PVSNet does not require the blending of sub-aperture views to generate the final synthesized view, setting it apart from the other methods in this comparison.
+Quantitative Processing Time: In terms of computational efficiency, PVSNet demonstrated significant improvements as shown in Table 4.
+Our PVSNet achieved an inference time of 0.0850 seconds for the whole LF reconstruction which included forward passes.
+However, PVSNet allows for direct querying of any target view within the light field baseline, enabling significantly faster per-view inference, averaging around 0.0079 seconds (126 fps).
+This efficiency highlights its suitability for real-time applications.
+In contrast, methods like LFSphereNet and DALF, despite achieving faster overall light field reconstruction times, necessitate the blending of sub-aperture views using per-view disk based blending algorithm (Overbeck et al., 2018).
+This process requires copying the light field data to the GPU buffer and utilizing shaders to blend the views for final synthesis.
+This additional step introduces an overhead of 1.301 seconds on an RTX 2070 Super, resulting in considerably slower time of 1.3018 seconds for LFSphereNet and 1.3587 seconds for DALF as shown in the last column of Table 4.
+Therefore, PVSNet presents a more efficient solution for narrow baseline view synthesis.
+| MLP | Normalization | Norm+PosEnc | Norm+MLP | Norm+PosEnc+MLP |
+| Positional Embedding | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
+| MLP | 23.6780 | 0.8267 | 0.8698 | 0.4577 | 0.1521 | 0.1063 |
+| Normalization | 25.5882 | 0.8466 | 0.9172 | 0.5673 | 0.1152 | 0.0766 |
+| Norm+Positional Encoding | 24.2862 | 0.8105 | 0.8719 | 0.4675 | 0.1376 | 0.0919 |
+| Norm+MLP | 26.7322 | 0.8581 | 0.9366 | 0.6422 | 0.1070 | 0.0888 |
+| Norm+PosEnc+MLP | 31.2258 | 0.9386 | 0.9793 | 0.8702 | 0.0613 | 0.0275 |
+Qualitative Image Quality: The qualitative evaluation of light field reconstruction is illustrated in Figure 7.
+We compare the visual performance of our proposed PVSNet against state-of-the-art methods for light field views, focusing on the top right corner sub-aperture view.
 The first two rows depict reconstruction results for the JPEG-Pleno (Rerabek and Ebrahimi, 2016) dataset.
-In these challenging scenarios, PLFNet demonstrates superior reconstruction quality with fewer visible artifacts, as evident in the error maps.
-Compared to LFSphereNet (Gond et al., 2023), which also shows strong performance, PLFNet exhibits more accurate details and reduced error magnitudes, highlighting its robustness and generalization capabilities.
+In these challenging scenarios, PVSNet demonstrates superior reconstruction quality with fewer visible artifacts, as evident in the error maps.
+Compared to LFSphereNet (Gond et al., 2023), which also shows strong performance, PVSNet exhibits more accurate details and reduced error magnitudes, highlighting its robustness and generalization capabilities.
 The subsequent rows display results for the Stanford Light Field Archive (Raj et al., 2016) dataset.
 Here, our method consistently outperforms NoisyLFRecon (Zhou et al., 2021), which utilizes nine input images for reconstruction.
-Despite this advantage, NoisyLFRecon produces noticeable artifacts and blurring in regions with fine structures.
-In contrast, PLFNet, which relies on only a single input image, achieves sharper reconstructions with more faithful reproduction of scene geometry and textures.
-This indicates that PLFNet can effectively capture and synthesize complex light field information even with minimal input data.
-Overall, the qualitative results corroborate the quantitative findings, showcasing ability of PLFNet to achieve high-quality reconstructions with low error rates.
+Despite this advantage in quantitative results where the average scores are higher, NoisyLFRecon produces noticeable artifacts and blurring in regions with fine structures.
+In contrast, PVSNet, which relies on only a single input image, achieves sharper reconstructions with more faithful reproduction of scene geometry and textures.
+This indicates that PVSNet can effectively capture and synthesize complex light field information even with minimal input data.
+More visual results can be found in our supplementary materials video, where we show the ability of PVSNet to generate any views within the LF baseline by directly querying the target position.
+Overall, the qualitative results corroborate the quantitative findings, showcasing ability of PVSNet to achieve high-quality reconstructions with low error rates.
 This performance shows potential for practical applications where high-fidelity light field reconstruction from sparse data is required.
@@
 In this section, we present a two-part ablation study to assess the effectiveness of our proposed method, focusing on positional understanding and the impact of different positional embedding techniques.
-The first part evaluates the performance of PLFNet+ under varying sparse input configurations, demonstrating how the number and location of input images influence view synthesis.
-The second part investigates the contribution of different positional encoding strategies to the overall reconstruction quality, highlighting the role of each component in our final model.
+The first part investigates the contribution of different positional embedding strategies to the overall reconstruction quality, highlighting the role of each component in our final model.
+The second part investigates the effect of modifying the structure of the Rendering Network, focusing on the contribution of Encoder I.
+Finally, the third part evaluates the performance of PVSNet under varying sparse input configurations, demonstrating how the number and location of input images influence view synthesis.
+5.1. Positional Embedding
+In the first part of the ablation study, we analyze the impact of different positional embedding strategies on view synthesis.
+We compare four configurations as shown in Figure 8: (i) MLP only, (ii) normalization only, (iii) normalization with positional encoding, (iv) normalization with MLP, and (v) the full setup with normalization, positional encoding, and MLP.
+This analysis aims to determine the contribution of each component in enhancing the model’s understanding of positional information.
+For a fair comparison, we train each variation of the network for same number of epochs.
+The results are summarized in Table 5.
+(i) MLP only: The model using only the MLP for the positional embedding block without any form of normalization or positional encoding exhibits the weakest performance.
+The lack of effective positional information results in inability of the rendering network to capture high-frequency variations in the scene, leading to a reconstruction that predominantly represents low-frequency features.
+This is reflected in a significantly lower average PSNR of 23.6780 dB, indicating poor view synthesis quality shown in Figure 9.
+(ii) Normalization only: When only coordinate normalization is employed, the network performs reasonably well for small baselines, where the scene geometry and appearance differences between the source and target viewpoints are minimal.
+However, as the baseline increases, the target viewpoint deviates more significantly from the source, leading to greater variations in scene appearance.
+Consequently, the model struggles to generalize to these larger baselines, yielding a PSNR of 25.5882 dB.
+This result highlights the limited capacity of normalization alone to effectively handle complex geometric transformations.
+(iii) Normalization with Positional Encoding: Introducing sinusoidal positional encoding on the normalized coordinates improves ability of the model to account for large geometric variations between source and target views.
+However, this approach also increases the search space of the rendering network as it must learn to map the given positional embedding vector to a suitable output space.
+This increased complexity can lead to suboptimal learning, as evidenced by the barbershop scene in Figure 9, where the rendered image exhibits a noticeable color mismatch.
+This configuration results in a PSNR of 24.2862 dB.
+(iv) Normalization with MLP: By combining normalization with an MLP, the model benefits from the ability of the MLP to project the normalized coordinates into a higher-dimensional space.
+This allows the rendering network to better capture both low and high-frequency variations in the scene, making it more effective at handling large baselines.
+The model achieves a PSNR of 26.7322 dB, demonstrating a marked improvement over the previous configurations.
+This result shows the importance of using an MLP to model complex transformations between source and target viewpoints, especially in scenarios involving significant changes in scene appearance.
+(v) Normalization, Positional Encoding, and MLP: The final configuration, which incorporates normalization, positional encoding, and an MLP, provides the most comprehensive solution.
+By applying positional encoding after normalization, the network benefits from a more structured and compact search space, allowing the MLP to more effectively map the positional inputs to a corresponding high dimensional space.
+This leads to a richer feature representation of the 3D positional information, facilitating the accurate reconstruction of complex novel views.
+As a result, this configuration achieves the highest PSNR of 31.2258 dB, demonstrating its superior performance in synthesizing novel views.
+Overall, the results from this ablation study highlight the significant contributions of each component in enhancing the understanding of positional information.
+The combination of normalization, positional encoding, and MLP provides the most robust and effective solution for accurate and high-quality view synthesis.
+This combination allows the model to better represent complex geometric and appearance changes, which is essential for synthesizing realistic novel views from a single input image.
+| Encoder I | Encoder II | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS | FPS |
+| ResNet | Ours | 31.2258 | 0.9386 | 0.9793 | 0.8702 | 0.0613 | 0.0275 | 62 |
+| DINOv2 | Ours | 32.5534 | 0.9529 | 0.9846 | 0.8945 | 0.0412 | 0.0258 | 23 |
+| - | Ours | 22.9666 | 0.8126 | 0.9288 | 0.6965 | 0.1251 | 0.0960 | 69 |
+| Number of Input Images | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
+| 2 Images - Random Location | 25.7561 | 0.8723 | 0.9255 | 0.6328 | 0.1021 | 0.0977 |
+| 1 Image - Random Location | 25.0533 | 0.8298 | 0.9041 | 0.5379 | 0.1227 | 0.1076 |
+| 1 Image - Central View | 31.2258 | 0.9386 | 0.9793 | 0.8702 | 0.0613 | 0.0275 |
+5.2. Rendering Network Components
+In the second part of the ablation study, we analyze the architectural choices made in the rendering network, particularly the use of a dual-encoder setup.
+The original architecture consists of two encoder branches: Encoder I: a ResNet-based feature extractor that processes the input image alone, and Encoder II: a custom designed encoder that jointly processes the input image and the positional embedding to generate position aware features.
+These features are later fused and passed through a decoder to produce the synthesized view.
+To understand the contribution of each encoder component, we evaluate three configurations (see Table 6): (i) the baseline setup using ResNet for Encoder I and our custom block for Encoder II, (ii) a variant where ResNet is replaced with DINOv2 (Oquab et al., 2023) to extract more semantically rich features, and (iii) a simplified setup where Encoder I is entirely removed, relying solely on Encoder II.
+The results demonstrate that while removing Encoder I simplifies the architecture and achieves the highest FPS (69), it significantly degrades synthesis quality across all metrics, including a PSNR drop of over 8 dB.
+This indicates that the positional features alone are insufficient for capturing fine appearance details without dedicated image encoding.
+This supports our architectural choice to include an additional image only encoder.
+Replacing ResNet with DINOv2 in Encoder I further improves quality, achieving the highest PSNR (32.55 dB), SSIM (0.95), and lowest perceptual error (LPIPS of 0.0258).
+However, this comes at a notable computational cost, reducing inference speed from 62 to 23 FPS.
+Given our real-time rendering objective, we select ResNet as the backbone for Encoder I to maintain a good balance between visual quality and speed.
+These findings validate the design choice of using a dual-encoder architecture, where Encoder I provides a strong appearance prior, and Encoder II injects pose specific spatial cues.
+The two branches complement each other, leading to sharper, more consistent novel views compared to relying solely on position-aware embeddings.
 5.3. Sparse Input Mode
-The first part of the ablation study explores how the performance of PLFNet+ is affected by varying the number and position of input images for view synthesis.
+The third part of the ablation study explores how the performance of PVSNet is affected by varying the number and position of input images for view synthesis.
 We evaluate three configurations: (i) a single central image as input, (ii) a single image from a random location within the cubical volume, and (iii) two images from random locations.
@@
 Table 7 presents the quantitative results of this study.
-As expected, the single central view input yields the best performance across all metrics, with a PSNR of 28.8023 and SSIM of 0.9009, indicating that the model performs best when provided with a centrally located reference image.
-The random input configurations, whether with one or two images, exhibit slightly reduced performance, suggesting that while the model can generalize to inputs from random locations, the quality of the synthesized view degrades with less structured input data.
-Nevertheless, even with these random inputs, PLFNet+ achieves acceptable performance, demonstrating its robustness in handling diverse input scenarios.
-| Number of Input Images | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
-| 2 Images - Random Location | 23.0571 | 0.7444 | 0.8462 | 0.4297 | 0.2782 | 0.2510 |
-| 1 Image - Random Location | 22.8429 | 0.7243 | 0.8386 | 0.3524 | 0.3373 | 0.3703 |
-| 1 Image - Central View | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 |
-6.2. Positional Embedding
-In the second part of the ablation study, we analyze the impact of different positional embedding strategies on view synthesis.
-We compare four configurations: (i) MLP only, (ii) normalization only, (iii) normalization with positional encoding, (iv) normalization with MLP, and (v) the full setup with normalization, positional encoding, and MLP.
-This analysis aims to determine the contribution of each component in enhancing the model’s understanding of positional information.
-The results are summarized in Table 6.
-The combination of normalization, positional encoding, and MLP consistently outperforms the other configurations, achieving a PSNR of 28.8023 and SSIM of 0.9003.
-This demonstrates the effectiveness of our final proposed method in capturing positional variations and synthesizing accurate novel views.
-By contrast, using only normalization or MLP leads to a noticeable drop in performance specially in high frequency regions, underscoring the importance of combining these techniques to achieve optimal results.
-The setup with normalization and MLP alone performs better than using MLP in isolation but still falls short of the complete configuration.
-These findings highlight the critical role of positional encoding and MLP in providing a richer representation of the 3D positional information, which is essential for accurately reconstructing complex novel views.
-| Positional Embedding | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
-| MLP | 23.2578 | 0.7633 | 0.8686 | 0.4063 | 0.2566 | 0.2726 |
-| Normalization | 25.2855 | 0.8376 | 0.9256 | 0.5309 | 0.2132 | 0.1680 |
-| Norm+Positional Encoding | 24.1945 | 0.7941 | 0.8962 | 0.4502 | 0.2276 | 0.1946 |
-| Norm+MLP | 26.4475 | 0.8508 | 0.9421 | 0.6247 | 0.2119 | 0.1647 |
-| Norm+PosEnc+MLP | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 |
-7. Conclusion
-In this paper we presented a novel, lightweight, position-aware network for real-time view synthesis and light field reconstruction, which leverages a position-aware embedding combined with a rendering network to achieve both efficiency and high visual quality.
-Through the development of PLFNet and its extension, PLFNet+, our approach demonstrates superior synthesis quality, particularly in handling complex translational movements without relying on computationally intensive operations like warping.
-Extensive evaluations reveal that PLFNet+ achieves a strong balance between visual fidelity and inference speed, positioning it as a feasible solution for applications requiring real-time performance, such as telepresence and augmented reality.
-Additionally, our network architecture shows robustness in various scenarios, from narrow to wide baseline settings, making it adaptable to dynamic and interactive environments.
-The promising results from ablation studies underscore the effectiveness of our positional embedding design, particularly the integration of normalization, positional encoding, and MLP components.
-Overall, our work provides a significant step towards more accessible, high-performance view synthesis frameworks, paving the way for broader application in consumer and industrial devices.
+As expected, the single central view input yields the best performance across all metrics, with a PSNR of 31.2258 dB and SSIM of 0.9386, indicating that the model performs best when provided with a centrally located reference image.
+The random input configurations, whether with one or two images, exhibit reduced performance, suggesting that while the model can generalize to inputs from random locations, the quality of the synthesized view degrades significantly with less structured input data.
+Nevertheless, even with these random inputs, PVSNet achieves acceptable performance, demonstrating its robustness in handling diverse input scenarios.
+6. Conclusion and Discussions
+This paper presents PVSNet and its efficient position-aware view synthesis framework, which offers a significant leap in the real-time generation of novel views from a single input image and a target position.
+Our method demonstrates superior efficiency, achieving real-time performance with a high inference speed, making it particularly suitable for interactive applications like immersive telepresence, 360∘ video with 6DoF and augmented reality (AR).
+Through extensive testing, PVSNet outperforms several state-of-the-art methods, such as AdaMPI (Han et al., 2022).
+Contribution to Existing Solutions: PVSNet enhances the field of view synthesis by introducing an effective solution for real-time inference.
+Compared to existing methods such as Quark (Flynn et al., 2024) or AdaMPI (Han et al., 2022), which require multi-view inputs or are computationally expensive, PVSNet excels in generating high-quality novel views from a single image.
+Its efficiency, particularly with large baselines, marks a significant improvement over previous approaches which require multiple images, or struggle with high computational demands or with generating continuous perspectives.
+While approaches like NViST (Jang and Agapito, 2024) and the use of Multi-Plane Images (MPI) have set the foundation for accurate scene representation, they often face limitations when applied in real-time applications due to their memory and computational overheads.
+PVSNet addresses these challenges by optimizing both computational efficiency and visual quality, thus supplementing existing solutions and opening up possibilities for their deployment in real-time industrial applications.
+Implications and Impact: The outcomes of this research hold substantial implications for both the academic community and industrial applications.
+For researchers, PVSNet provides a more accessible approach to view synthesis, offering a platform for further innovation in the area of real-time view synthesis.
+The use of positional embeddings for view synthesis could inspire further exploration of more efficient representations using smaller vision transformer models for real-time scene understanding.
+For engineers and industrial applications, the real-time nature of PVSNet makes it a powerful tool for sectors such as remote operations, training simulations, and VR experiences.
+By providing continuous perspective views with minimal latency, this framework could drastically improve user interactions in these fields.
+Furthermore, the ability to synthesize novel views from a single input image opens up new possibilities for systems with limited computational resources, where traditional multi-view methods might not be feasible.
+However, while the method achieves impressive results, there are still areas for improvement.
+One limitation is the handling of occlusions and complex geometries in highly dynamic environments when the rendering volume is large for novel view synthesis.
+Future Directions: As PVSNet serves as a foundation for real-time position-aware view synthesis, future work can extend this approach by exploring the way to use the latent space generated by the model for other scene understanding tasks, such as depth estimation, segmentation and object detection, to further enhance its applicability in diverse environments.
+Future research could explore ways to further enhance depth estimation or integrate multi-view information in a hybrid manner, balancing computational efficiency with the ability to handle more complex scene interactions.
+Additionally, expanding its generalization capabilities to handle a wider range of real-world scenes and augmenting its robustness for large-scale deployment will be crucial steps toward achieving broader industry adoption.
 Acknowledgements.
 The work was supported by the European Joint Doctoral Programme on Plenoptic Imaging (PLENOPTIMA) through the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No.
-956770.
+956770, by the Knowledge Foundation, Sweden, with grant number 2019-0251, and by Mid Sweden University internal funding.
 The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no.
@@
 - (1)
-- Bae et al.
-(2021) Kyuho Bae, Andre Ivan, Hajime Nagahara, and In Kyu Park.
-2021.
-5d light field synthesis from a monocular video.
-In 2020 25th International Conference on Pattern Recognition (ICPR).
-IEEE, 7157–7164.
+- Ahn and Kim (2013) Ilkoo Ahn and Changick Kim. 2013. A novel depth-based virtual view synthesis method for free viewpoint video. IEEE Transactions on Broadcasting 59, 4 (2013), 614–626.
 - Barron et al.
@@
 5855–5864.
+- Bertel et al.
+(2019) Tobias Bertel, Neill DF Campbell, and Christian Richardt.
+2019.
+Megaparallax: Casual 360 panoramas with motion parallax.
+IEEE transactions on visualization and computer graphics 25, 5 (2019), 1828–1835.
 - Caesar et al.
@@
 1209–1218.
-- Chandramouli et al.
-(2020) Paramanand Chandramouli, Kanchana Vaishnavi Gandikota, Andreas Goerlitz, Andreas Kolb, and Michael Moeller.
-2020.
-A generative model for generic light field reconstruction.
-IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 4 (2020), 1712–1724.
-- Chen et al.
-(2020) Bin Chen, Lingyan Ruan, and Miu-Ling Lam.
-2020.
-LFGAN: 4D Light Field Synthesis from a Single RGB Image.
-ACM Trans.
-Multimedia Comput.
-Commun.
-Appl.
-16 (2 2020).
-Issue 1.
-https://doi.org/10.1145/3366371
+- Chandan et al.
+(2021) Kishan Chandan, Jack Albertson, Xiaohan Zhang, Xiaoyang Zhang, Yao Liu, and Shiqi Zhang.
+2021.
+Learning to guide human attention on mobile telepresence robots with 360 vision.
+In 2021 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS).
+IEEE, 5297–5304.
+- Chen and Koltun (2017) Qifeng Chen and Vladlen Koltun.
+2017.
+Photographic image synthesis with cascaded refinement networks.
+In Proceedings of the IEEE international conference on computer vision.
+1511–1520.
+- Chen and Williams (1993) Shenchang Eric Chen and Lance Williams.
+1993.
+View interpolation for image synthesis.
+In Proceedings of the 20th Annual Conference on Computer Graphics and Interactive Techniques (Anaheim, CA) (SIGGRAPH ’93).
+Association for Computing Machinery, New York, NY, USA, 279–288.
+https://doi.org/10.1145/166117.166153
 - Cun et al.
@@
 13142–13153.
+- Dima et al.
+(2020) Elijs Dima, Kjell Brunnström, Mårten Sjöström, Mattias Andersson, Joakim Edlund, Mathias Johanson, and Tahir Qureshi.
+2020.
+Joint effects of depth-aiding augmentations and viewing positions on the quality of experience in augmented telepresence.
+Quality and User Experience 5 (2020), 1–17.
 - Dima and Sjöström (2021) Elijs Dima and Mårten Sjöström. 2021. Camera and Lidar-Based View Generation for Augmented Remote Operation in Mining Applications. IEEE Access 9 (2021), 82199–82212.
@@
 - Dosovitskiy (2020) Alexey Dosovitskiy. 2020. An image is worth 16x16 words: Transformers for image recognition at scale. arXiv preprint arXiv:2010.11929 (2020).
+- Fehn (2004) Christoph Fehn.
+2004.
+Depth-image-based rendering (DIBR), compression, and transmission for a new approach on 3D-TV.
+In Stereoscopic displays and virtual reality systems XI, Vol.
+5291.
+SPIE, 93–104.
 - Flynn et al.
@@
 In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR).
+- Flynn et al.
+(2024) John Flynn, Michael Broxton, Lukas Murmann, Lucy Chai, Matthew DuVall, Clément Godard, Kathryn Heal, Srinivas Kaza, Stephen Lombardi, Xuan Luo, Supreeth Achar, Kira Prabhu, Tiancheng Sun, Lynn Tsai, and Ryan Overbeck.
+2024.
+Quark: Real-time, High-resolution, and General Neural View Synthesis.
+ACM Trans.
+Graph.
+43, 6, Article 194 (Nov.
+2024), 20 pages.
+https://doi.org/10.1145/3687953
 - Fridovich-Keil et al.
@@
 297–306.
-- Ivan et al. (2019) Andre Ivan, In Kyu Park, et al. 2019. Synthesizing a 4D spatio-angular consistent light field from a single image. arXiv preprint arXiv:1903.12364 (2019).
+- Jampani et al.
+(2021) Varun Jampani, Huiwen Chang, Kyle Sargent, Abhishek Kar, Richard Tucker, Michael Krainin, Dominik Kaeser, William T Freeman, David Salesin, Brian Curless, et al.
+2021.
+Slide: Single image 3d photography with soft layering and depth-aware inpainting.
+In Proceedings of the IEEE/CVF International Conference on Computer Vision.
+12518–12527.
 - Jang and Agapito (2024) Wonbong Jang and Lourdes Agapito.
@@
 1169–1178.
+- Konrad et al.
+(2017) Robert Konrad, Donald G Dansereau, Aniq Masood, and Gordon Wetzstein.
+2017.
+Spinvr: towards live-streaming 3d virtual reality video.
+ACM Transactions on Graphics (TOG) 36, 6 (2017), 1–12.
 - Levoy and Hanrahan (1996) Marc Levoy and Pat Hanrahan. 1996. Light field rendering. In Proceedings of the 23rd annual conference on Computer graphics and interactive techniques. 31–42.
@@
 ACM transactions on graphics (TOG) 41, 4 (2022), 1–15.
+- Oquab et al.
+(2023) Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, et al.
+2023.
+Dinov2: Learning robust visual features without supervision.
+arXiv preprint arXiv:2304.07193 (2023).
 - Overbeck et al.
@@
 - Raj et al. (2016) Abhilash Sunder Raj, Michael Lowney, Raj Shah, and Gordon Wetzstein. 2016. Stanford lytro light field archive.
+- Ranftl et al.
+(2022) René Ranftl, Katrin Lasinger, David Hafner, Konrad Schindler, and Vladlen Koltun.
+2022.
+Towards Robust Monocular Depth Estimation: Mixing Datasets for Zero-Shot Cross-Dataset Transfer.
+IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 3 (2022), 1623–1637.
+https://doi.org/10.1109/TPAMI.2020.3019967
 - Rerabek and Ebrahimi (2016) Martin Rerabek and Touradj Ebrahimi. 2016. New light field image dataset. In 8th International Conference on Quality of Multimedia Experience (QoMEX).
+- Sakashita et al.
+(2023) Mose Sakashita, Hyunju Kim, Brandon Woodard, Ruidong Zhang, and François Guimbretière.
+2023.
+VRoxy: Wide-Area Collaboration From an Office Using a VR-Driven Robotic Proxy.
+In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (San Francisco, CA, USA) (UIST ’23).
+Association for Computing Machinery, New York, NY, USA, Article 59, 13 pages.
+https://doi.org/10.1145/3586183.3606743
 - Seitz and Dyer (1999) Steven M Seitz and Charles R Dyer. 1999. Photorealistic scene reconstruction by voxel coloring. International journal of computer vision 35 (1999), 151–173.
 - Sheikh and Bovik (2006) Hamid R Sheikh and Alan C Bovik. 2006. Image information and visual quality. IEEE Transactions on image processing 15, 2 (2006), 430–444.
+- Shih et al.
+(2020) Meng-Li Shih, Shih-Yang Su, Johannes Kopf, and Jia-Bin Huang.
+2020.
+3d photography using context-aware layered depth inpainting.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+8028–8038.
+- Solovev et al.
+(2023) Pavel Solovev, Taras Khakhulin, and Denis Korzhenkov.
+2023.
+Self-improving multiplane-to-layer images for novel view synthesis.
+In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision.
+4309–4318.
 - Srinivasan et al.
@@
 https://openaccess.thecvf.com/content_iccv_2017/html/Srinivasan_Learning_to_Synthesize_ICCV_2017_paper.html
+- Suhail et al.
+(2022) Mohammed Suhail, Carlos Esteves, Leonid Sigal, and Ameesh Makadia.
+2022.
+Generalizable patch-based neural rendering.
+In European Conference on Computer Vision.
+Springer, 156–174.
 - Tucker and Snavely (2020) Richard Tucker and Noah Snavely.
@@
 551–560.
+- Waidhofer et al.
+(2022) John Waidhofer, Richa Gadgil, Anthony Dickson, Stefanie Zollmann, and Jonathan Ventura.
+2022.
+PanoSynthVR: Toward Light-weight 360-Degree View Synthesis from a Single Panoramic Input.
+In 2022 IEEE International Symposium on Mixed and Augmented Reality (ISMAR).
+IEEE, 584–592.
 - Wang et al.
@@
 arXiv preprint arXiv:2210.04628 (2022).
+- Yao et al. (2018) Yao Yao, Zixin Luo, Shiwei Li, Tian Fang, and Long Quan. 2018. MVSNet: Depth Inference for Unstructured Multi-view Stereo. European Conference on Computer Vision (ECCV) (2018).
 - Yu et al.
@@
 https://doi.org/10.1109/QoMEX61742.2024.10598264
+- Zhang et al.
+(2024) Jason Y Zhang, Amy Lin, Moneish Kumar, Tzu-Hsuan Yang, Deva Ramanan, and Shubham Tulsiani.
+2024.
+Cameras as rays: Pose estimation via ray diffusion.
+arXiv preprint arXiv:2402.14817 (2024).
 - Zhang et al. (2018) Richard Zhang, Phillip Isola, Alexei A Efros, Eli Shechtman, and Oliver Wang. 2018. The Unreasonable Effectiveness of Deep Features as a Perceptual Metric. In CVPR.
@@
 https://doi.org/10.1016/j.sigpro.2021.108121
+- Zhu et al. (2019) Shiping Zhu, Hao Xu, and Lina Yan. 2019. An improved depth image based virtual view synthesis method for interactive 3D video. IEEE Access 7 (2019), 115171–115180.